### v1

# Flat quasi-Frobenius Lie superalgebras

Abstract.
A non-associative superalgebra is called pre-symplectic if it is equipped with a non-degenerate, anti-symmetric bilinear form. It is called quasi-Frobenius if, in addition, is a Lie superalgebra and the form is closed. We introduce the Levi-Civita product associated with pre-symplectic superalgebras and establish its existence and uniqueness. We then introduce the symplectic product associated with quasi-Frobenius Lie superalgebras. We prove that while such a product always exists, it is not unique. We therefore define a natural symplectic product that depends only on the Lie structure and the bilinear form. When the curvature of this product vanishes, the superalgebra is called a flat quasi-Frobenius Lie superalgebra.
In this paper, we study flat quasi-Frobenius Lie superalgebras and introduce the notion of a flat double extension. We prove that the double extension process characterizes such superalgebras. More precisely, every flat orthosymplectic (resp. periplectic) quasi-Frobenius Lie superalgebra can be obtained by a sequence of flat double extensions starting from an abelian one (resp. the trivial one). Moreover, we show that every flat quasi-Frobenius Lie superalgebra is nilpotent with a degenerate center. We apply our results to obtain a complete classification of such superalgebras of total dimension at most five.
Key words and phrases:
Quasi-Frobenius Lie (super)algebras, the Levi-Civita product, symplectic product, double extension.
2020 Mathematics Subject Classification:
17A70; 17B05; 17A60; 17D25
S.B. was supported by the grant NYUAD-065
1. Introduction
1.1. The Riemannian setting
A pseudo-Riemannian Lie group is a Lie group $G$ endowed with a left-invariant pseudo-Riemannian metric $\mu$. Its Lie algebra $\mathfrak{g}=T_{e}G$, equipped with the scalar product $\langle\;,\;\rangle=\mu_{e}$, is called a pseudo-Riemannian Lie algebra. Whenever there exists a torsion-free connection compatible with the metric, such a connection is necessarily unique and is called the Levi-Civita connection (see for instance [NOM, ON]). The Levi-Civita connection induces a bilinear product $(u,v)\mapsto u\star v$ on $\mathfrak{g}$, called the Levi-Civita product, defined by Koszul’s formula:
$2\langle u\star v,w\rangle=\langle[u,v],w\rangle+\langle[w,u],v\rangle+\langle[w,v],u\rangle,\qquad\text{for all }u,v,w\in\mathfrak{g}.$
This product satisfies the following properties:
$u\star v-v\star u=[u,v],\qquad\langle u\star v,w\rangle=-\,\langle v,u\star w\rangle,\qquad\text{for all }u,v,w\in\mathfrak{g}.$
In the setting of pseudo-Euclidean non-associative superalgebras (in particular, Leibniz superalgebras) this construction has been generalized in [BBE2].
The curvature of $(\mathfrak{g},\langle\;,\;\rangle)$ at the identity is
$K(u,v):={\mathrm{L}}_{[u,v]}-[{\mathrm{L}}_{u},{\mathrm{L}}_{v}],\qquad\text{for all }u,v\in\mathfrak{g},$
where ${\mathrm{L}}_{u}$ is the left multiplication defined by ${\mathrm{L}}_{u}(v)=u\star v,$ for all $u,v\in\mathfrak{g}.$
A pseudo-Riemannian Lie group $(G,\mu)$ is called flat when its curvature vanishes identically; in this case, $(\mathfrak{g},\langle\;,\;\rangle)$ will be refered to as a flat pseudo-Riemannian Lie algebra. In fact, flatness is equivalent to $(\mathfrak{g},\star)$ being a left-symmetric algebra..
The study of flat Lie algebras originated with the Euclidean case. A foundational result was proved by Milnor in 1976 [M], who characterized such algebras: a Lie algebra $\mathfrak{g}$ is flat Euclidean if and only if it admits an orthogonal decomposition $\mathfrak{g}=\mathfrak{b}\oplus\mathfrak{u}$ where $\mathfrak{u}$ is an abelian ideal, $\mathfrak{b}$ is an abelian subalgebra, and for every $b\in\mathfrak{b}$, the adjoint map $\mathop{\mathrm{ad}}\nolimits_{\mathfrak{b}}$ is skew-symmetric with respect to the inner product.
In recent years, a number of papers have investigated flat pseudo-Riemannian Lie algebras with various signatures, using in particular the method of double extension introduced in [AM]. Further developments include the study of flat Lorentzian Lie algebras [AM, ABL, BL] as well as the study of flat pseudo-Euclidean nilpotent Lie algebras with the signature (2,n2)(2,n-2), see [BL1]. Furthermore, in [BBE2], the authors studied flat pseudo-Euclidean left-Leibniz superalgebras.
1.2. The symplectic setting.
A symplectic Lie group is a Lie group $G$ endowed with a left-invariant symplectic form $\Omega$. The associated Lie algebra $\mathfrak{g}=T_{e}G$, equipped with $\omega=\Omega_{e}$, is a symplectic Lie algebra (called quasi-Frobenuis in this paper). A connection $\nabla$ on $G$ is called a symplectic connection if it is torsion-free and is compatible with $\Omega$, i.e., $\nabla\Omega=0$. Such a symplectic connection always exists, but it is not unique (see [BCGRS]). Whenever a symplectic connection is given, one may define a bilinear product $(u,v)\mapsto u\bullet v$ on $\mathfrak{g}$ that satisfies the following properties:
$u\bullet v-v\bullet u=[u,v],\qquad\omega(u\bullet v,w)=-\,\omega(v,u\bullet w),\qquad\text{for all }u,v,w\in\mathfrak{g}.$
In [BB], the authors introduced a natural symplectic connection on symplectic Lie groups, defined by
$\omega(u\star v,w)=\tfrac{1}{3}\,\omega([u,v],w)+\tfrac{1}{3}\,\omega([u,w],v),\qquad\text{for all }u,v,w\in\mathfrak{g}.$
As in the pseudo-Riemannian case, one defines the curvature of this connection. A symplectic Lie group $(G,\Omega)$ is called flat when its curvature vanishes identically. In this case, the induced Lie algebra $(\mathfrak{g},\omega)$ is a flat quasi-Frobenius Lie algebra. In fact, flatness is equivalent to the fact that the symplectic product defines a left-symmetric structure on $\mathfrak{g}$.
Despite its importance, the flat symplectic case has received little attention from experts. General results on flat quasi-Frobenius Lie algebras appear in [V]. Meanwhile, the symplectic connection above was studied in [BELL]; further contributions can be found in [A, BC, MR1, NB].
The study of such structures is motivated by a well-known open problem of Milnor [M], which seeks a classification and description of flat affine Lie groups, i.e. Lie groups carrying a left-invariant flat affine structure.
Recall that if $(G,\Omega)$ is a symplectic Lie group, then the left-invariant connection $\nabla$ defined by
(1) $\omega(\nabla_{u}v,w)=-\omega(v,[u,w]),\qquad\text{for all }u,v,w\in\mathfrak{g},$
is indeed flat, i.e. its curvature vanishes. Note, in addition, that $\nabla$ is symplectic (that is, $\nabla\Omega=0$) if and only if $G$ is abelian, equivalently its Lie algebra $\mathfrak{g}$ is abelian (see [Ch]).
1.3. Outline of the paper
Section 2 collects the preliminary material required for our work. For completeness and to establish notation, we include a review of key concepts, some of which may be familiar to specialists.
In Section 3, we introduce the Levi-Civita product on a class of non-associative anti-commutative superalgebras that we call pre-symplectic. These superalgebras admit a non-degenerate anti-symmetric bilinear form $\omega$ (not necessarily closed). This product $\bullet$ is the unique bilinear map on the pre-symplectic superalgebra $\mathcal{A}$} that satisfies the following properties:
$u,v]=u\bullet v-(-1)^{|u||v|}v\bullet u,\qquad\omega(u\bullet v,w)=(-1)^{|u||v|}\,\omega(v,u\bullet w),$
for all homogeneous $u,v,w\in{\mathcal{A}}$. The product $\bullet$ is also characterized by a Koszul formula (see Eq. 4). Its existence and uniqueness are established in Prop 3.1.1.  Proposition.
Further, we study the Levi-Civita product in the case where the pre-symplectic superalgebra is a Lie superalgebra and the form $\omega$ is closed. We call these superalgebras quasi-Frobenius Lie superalgebras, following [BBE, BE, BM]. It is worth mentioning that several authors call them symplectic Lie superalgebras (see [BC, BEB, BBNS] to cite a few). We show that the Levi-Civita product associated with a quasi-Frobenius Lie superalgebra $(\mathfrak{g},\omega)$ is always right-symmetric (see Prop. 3.1.5.  Proposition) and satisfies
$\omega(u\bullet v,w)=\omega(u,[v,w]),\qquad\text{ for all homogeneous }\,u,v,w\in\mathfrak{g}.$
The opposite product given by $u\circ v=-(-1)^{|u||v|}v\bullet u,$ for all homogeneous $u,v\in\mathfrak{g}$, defines a left-symmetric structure on $\mathfrak{g}$}, which superize the classical construction in Eq.(1).
In addition, we introduce a second product $\star$ on a quasi-Frobenius Lie superalgebra $(\mathfrak{g},\omega)$, defined by
$u\star v-(-1)^{|u||v|}v\star u=[u,v],\qquad\omega(u\star v,w)=-(-1)^{|v||w|}\,\omega(v,u\star w).$
Such a symplectic product exists but is not unique (see Corollary 3.2.8.  Corollary). This contrasts with the uniqueness of the Levi-Civita product on pre-symplectic superalgebras as well as in the case of pseudo-Euclidean superalgebras (see [BBE2]).
We show nevertheless that in any quasi-Frobenius Lie superalgebra $(\mathfrak{g},\omega)$ we can construct a unique symplectic product with the key property that it depends only on the Lie bracket and the symplectic form $\omega$. This product is explicitly given by:
$\omega(u\star v,w)=\frac{1}{3}\left(\omega([u,v],w)+(-1)^{|v||w|}\,\omega([u,w],v)\right),\text{ for all homogeneous }\,u,v,w\in\mathfrak{g}.$
We call this product the natural symplectic product associated with $(\mathfrak{g},[\;,\;],\omega)$. When the curvature of the natural symplectic product $\star$ is zero, the product is called flat, and the triple $(\mathfrak{g},[\;,\;],\omega)$ is referred to as a flat quasi-Frobenius Lie superalgebra. In fact, flatness is equivalent to the fact that the product defines a left-symmetric structure on the underlying vector space of $\mathfrak{g}$}.
In Section 4, we initiate the study of the structure theory for flat quasi-Frobenius Lie superalgebras. We show that such algebras are solvable when the form $\omega$ is even (Theorem 4.4.  Theorem) and nilpotent when the form $\omega$ is odd (Theorem 4.5.  Theorem).
In Section 5, we introduce the notion of flat double extension for flat quasi-Frobenius Lie superalgebras, see Theorem 5.5.5.  Theorem, Theorem 5.5.12.  Theorem, Theorem 5.5.15.  Theorem, and Theorem 5.5.19.  Theorem. This construction is inspired by the classical double extension, originally introduced by Medina and Revoy [MR] for quadratic Lie algebras, and its symplectic analogs [MR2, DM]. It is worth noticing that this notion differs from the notion of double extensions of quasi-Frobenius Lie superalgebras considered in [BM, BE].
In Section 6, we show that all non-abelian flat quasi-Frobenius Lie superalgebras have a degenerate center, see Theorem 6.1.  Theorem. Moreover, every such Lie superalgebra can be obtained via flat double extensions. In particular, every flat orthosymplectic quasi-Frobenius Lie superalgebra can be constructed by a sequence of flat double extensions starting from an an abelian Lie superalgebra (see Corollary 6.8.  Corollary), while flat periplectic quasi-Frobenius Lie superalgebras can be obtained by a sequence of flat double extensions starting from the trivial algebra $\{0\}$ (see Corollary 6.9.2.  Corollary). Furthermore, we show that every flat orthosymplectic quasi-Frobenius Lie superalgebra is not only solvable but nilpotent (Theorem 6.9.  Theorem).
Section 7 is devoted to classifying all flat quasi-Frobenius Lie superalgebras of total dimension at most five over an algebraically closed field of characteristic zero.
2. Backgrounds
Let $\mathbb{K}$ be an arbitrary field of characteristic zero. The group of integers modulo $2$ is denoted by $\mathbb{Z}_{2}$.
Let $V=V_{\bar{0}}\oplus V_{\bar{1}}$ be a superspace defined over $\mathbb{K}$}. The parity of a homogeneous element $v\in V_{\bar{i}}$ is denoted by $|v|:=\bar{i}$. The element $v$ is called even if $v\in V_{\bar{0}}$ and odd if $v\in V_{\bar{1}}$.
Throughout the text, all elements in linear expressions are assumed to be homogeneous, unless otherwise stated.
Let $V$ be a vector superspace. For homogeneous elements $x,y\in V$ we denote by $x^{*},y^{*}\in V^{*}$ their duals . For every $x^{*}\otimes y^{*}\in V^{*}\otimes V^{*}$ and $u\otimes v\in V\otimes V$, we define the pairing
$\langle x^{*}\otimes y^{*},u\otimes v\rangle:=(-1)^{|y||u|}x^{*}(u)\,y^{*}(v).$
The wedge (graded antisymmetrisation) is given by
$x^{*}\wedge y^{*}=x^{*}\otimes y^{*}-(-1)^{|x||y|}\,y^{*}\otimes x^{*}.$
A linear map $\varphi:V\rightarrow W$ between two superspaces is called even if $\varphi(V_{\bar{i}})\subset W_{\bar{i}}$ and odd if $\varphi(V_{\bar{i}})\subset W_{\overline{i}+\overline{1}}$.
Let $V$ and $W$ be two superspaces. Let ${\mathcal{B}}\in\text{Bil}(V,W)$ be a homogeneous bilinear form. Recall that the Gram matrix $B=(B_{ij})$ associated to $\mathcal{B}$ is given by the following formula:
[…]
This definition allows us to identify a bilinear form $B(V,W)$ with an element of $\mathrm{Hom}(V,W^{*})$. Consider the upsetting of bilinear forms $u:\text{Bil}(V,W)\rightarrow\text{Bil}(W,V)$ given by the formula
[…]
In terms of the Gram matrix $B$ of $\mathcal{B}$}, we have
$u(B)=\left(\begin{array}[]{cc}R^{t}&(-1)^{p(B)}T^{t}\\ (-1)^{p(B)}S^{t}&-U^{t}\end{array}\right),\text{ for $B=\left(\begin{array}[]{cc}R&S\\ T&U\end{array}\right)$}.$
From now on, we will assume that $V=W$. Following [L], we say that:
  • the form $\mathcal{B}$ is symmetric if and only if $u(B)=B$;
  • the form $\mathcal{B}$ is anti-symmetric if and only if u(B)=Bu(B)=-B.
Given a linear map $f:V\rightarrow V$. The adjoint of $f$ with respect to $\mathcal{B}$} is the linear map $f^{*}$ defined as follows:
${\mathcal{B}}(f(v),w)=(-1)^{|f||v|}{\mathcal{B}}(v,f^{*}(w)).$
We say that $f$ is $\mathcal{B}$}-symmetric (resp. $\mathcal{B}$}-antisymmetric) if $f^{*}=f$ (resp. $f^{*}=-f$).
A sub-superspace $U\subseteq V$ is called degenerate if and only if $U\cap U^{\perp}\not=\{0\}$, where $\perp$ stands for the orthogonal with respect to the bilinear form $\mathcal{B}$}.
Throughout this paper, $\omega$ will denote an anti-symmetric bilinear form, which by definition satisfies
$\omega(u,v)=-(-1)^{|u||v|}\omega(v,u)\text{ for every }u,v\in V.$
A non-degenerate, anti-symmetric even form is called orthosymplectic; an odd one is called periplectic.
Lemma (See [BM]).
Let $V=V_{\bar{0}}\oplus V_{\bar{1}}$ be a finite-dimensional superspace with $V_{\bar{1}}\neq\{0\}$, endowed with a periplectic form $\omega$. Then, $\dim(V_{\bar{0}})=\dim(V_{\bar{1}})$.
A superalgebra $({\mathcal{A}},\bullet)$ over a field $\mathbb{K}$ consists of a superspace ${\mathcal{A}}={\mathcal{A}}_{\bar{0}}\oplus{\mathcal{A}}_{\bar{1}}$ and a binary operation satisfying: ${\mathcal{A}}_{\alpha}\bullet{\mathcal{A}}_{\beta}\subseteq{\mathcal{A}}_{\alpha+\beta},\forall\alpha,\beta\in\mathbb{Z}_{2}$.
For every $u\in{\mathcal{A}}$, we denote the left and right multiplication operators by ${\mathrm{L}}_{u}$ and ${\mathrm{R}}_{u}$, respectively. More precisely, ${\mathrm{L}}_{u}$ and ${\mathrm{R}}_{u}$ are defined as follows:
${\mathrm{L}}^{\bullet}_{u}(v):=u\bullet v\text{ and }{\mathrm{R}}^{\bullet}_{u}(v):=(-1)^{|u||v|}v\bullet u,\text{ for all $v\in{\mathcal{A}}$}.$
Let us define:
$N_{\ell}(\mathfrak{g},\bullet):=\left\{u\in\mathfrak{g},\mathrm{L}^{\bullet}_{u}=0\right\},\;N_{r}(\mathfrak{g},\bullet):=\left\{u\in\mathfrak{g},\mathrm{R}_{u}^{\bullet}=0\right\},\;\text{and}\;N(\mathfrak{g},\bullet):=N_{\ell}(\mathfrak{g},\bullet)\cap N_{r}(\mathfrak{g},\bullet).$
We refer to these as the left normalizer, right normalizer, and (two-sided) normalizer, respectively.
Let $({\mathcal{A}},\bullet)$ be a non-associative superalgebra. On the underlying vector superspace $\mathcal{A}$}, we define the corresponding commutator
$u,v]_{\bullet}:=u\bullet v-(-1)^{|u||v|}v\bullet u,\;u,v\in{\mathcal{A}}.$
The algebra $({\mathcal{A}},[\;,\;]_{\bullet})$ will be denote by $\mathcal{A}$$^{-}$. The algebra $({\mathcal{A}},\bullet)$ is called Lie-admissible superalgebra, if $\mathcal{A}$$^{-}$ is a Lie superalgebra. In this case, $\mathcal{A}$$^{-}$ is called the sub-adjacent Lie superalgebra of $({\mathcal{A}},\bullet)$ and the latter is called a compatible (Lie-admissible) superalgebra structure on the Lie superalgebra $\mathcal{A}$$^{-}$.
A superalgebra $({\mathcal{A}},\bullet)$ is called left-symmetric (resp. right-symmetric) if and only if
$\mathrm{as}(a,b,c)=(-1)^{|a||b|}\,\mathrm{as}(b,a,c),\quad(\text{resp. }\mathrm{as}(a,b,c)=(-1)^{|b||c|}\,\mathrm{as}(a,c,b)),$
where $\mathrm{as}(a,b,c)=(a\bullet b)\bullet c-a\bullet(b\bullet c)$ is the associator.
It is well-known that such superalgebras are Lie-admissible.
A pre-symplectic superalgebra $({\mathcal{A}},\cdot,\omega)$ is a non-associative superalgebra $({\mathcal{A}},\cdot)$ endowed with a non-degenerate anti-symmetric bilinear form $\omega$.
3. The Levi-Civita product and the Symplectic product on quasi-Frobenius Lie superalgebras
3.1. The Levi-Civita product
In [BBE2], the authors introduced the Levi-Civita product associated with a non-associative superalgebra equipped with a non-degenerate symmetric bilinear form. Here, we introduce its analogue in the case where the form is anti-symmetric.
Proposition.
Let $({\mathcal{A}},[\;,\;],\omega)$ be a pre-symplectic anti-commutative superalgebra. There exists a unique product $\bullet$ on $\mathcal{A}$} that satisfies (for all $u,v,w\in{\mathcal{A}})$:
(2) $[u,v]_{\bullet}$ $=$ $[u,v],$
(3) $(\mathrm{L}_{u}^{\bullet})^{*}$ $=$ $\mathrm{L}_{u}^{\bullet}.$
More precisely, the product $\bullet$ is defined by
(4) $2\omega(u\bullet v,w)=\omega([u,v],w)-(-1)^{|u||v|+|w||u|}\omega([v,w],u)+(-1)^{|v||w|+|u||w|}\omega([w,u],v).$
The product $\bullet$ is called the Levi-Civita product associated with $({\mathcal{A}},[\;,\;],\omega)$.
Proof.
(Uniqueness) Suppose that there exist two products $\bullet$ and $\bullet^{\prime}$ satisfying conditions (2), (3). For all $u,v,w\in\mathfrak{g}$, we have
$\omega(u\bullet v,w)$ $=\omega([u,v],w)+(-1)^{|u||v|}\omega(v\bullet u,w)$
$=\omega([u,v],w)-(-1)^{|u||v|+|w||u|}\omega(v\bullet w,u)$
$=\omega([u,v],w)-(-1)^{|u||v|+|w||u|}\omega([v,w],u)-(-1)^{|u||v|+|w||u|+|v||w|}\omega(w\bullet v,u)$
$=\omega([u,v],w)-(-1)^{|u||v|+|w||u|}\omega([v,w],u)+(-1)^{|v||w|+|u||w|}\omega(w\bullet u,v)$
$=\omega([u,v],w)-(-1)^{|u||v|+|w||u|}\omega([v,w],u)+(-1)^{|v||w|+|u||w|}\omega([w,u],v)-\omega(u\bullet v,w).$
The same computation by means of $\bullet^{\prime}$, implies that $2\,\omega(u\bullet v,w)=2\,\omega(u\bullet^{\prime}v,w).$ Since $\omega$ is non-degenerate, it follows that $\bullet=\bullet^{\prime}$.
(Existence) We define the product $\bullet$ by the following relation:
$2\,\omega(u\bullet v,w)=\omega([u,v],w)-(-1)^{|u||v|+|w||u|}\omega([v,w],u)+(-1)^{|v||w|+|u||w|}\omega([w,u],v),$
for all $u,v,w\in{\mathcal{A}}$. It is straightforward to verify that this product satisfies
$\omega(\mathrm{L}_{u}^{\bullet}(v),w)=-(-1)^{|u||v|}\omega(v,\mathrm{L}_{u}^{\bullet}(w)),\text{ and }[u,v]_{\bullet}=[u,v].$
Hence, the product $\bullet$ exists and is unique. ∎
Let $(\mathfrak{g},[\;,\;],\omega)$ be a pre-symplectic Lie superalgebra, and let $\bullet$ denote its Levi-Civita product.
Eqns. (2) and (3) imply that the multiplication operator $\mathrm{L}^{\bullet}_{u}$, is $\omega$-symmetric, for every $u\in\mathfrak{g}$, and that $\mathop{\mathrm{ad}}\nolimits_{u}=\mathrm{L}^{\bullet}_{u}-\mathrm{R}^{\bullet}_{u}$, where $\mathop{\mathrm{ad}}\nolimits_{u}:\mathfrak{g}\to\mathfrak{g}$ is the adjoint operator. Let us denote by $Z(\mathfrak{g})$ the center of $\mathfrak{g}$}. The following relations hold (where $\mathfrak{g}\bullet\mathfrak{g}={u\bullet v\mid u,v\in\mathfrak{g}}$):
[…]
Definition.
Let $(\mathfrak{g},\cdot,\omega)$ be a pre-symplectic superalgebra. The form $\omega$ is said to be closed if, for all elements $u,v,w\in\mathfrak{g}$, the following identity holds:
(6) $(-1)^{|u||w|}\omega(u\cdot v,w)+(-1)^{|v||u|}\omega(v\cdot w,u)+(-1)^{|w||v|}\omega(w\cdot u,v)=0.$
Following [BBE, BE, BM], if a pre-symplectic Lie superalgebra $(\mathfrak{g},[\;,\;],\omega)$ satisfies the condition (6), then it is called a quasi-Frobenius Lie superalgebra. Cohomologically speaking, this condition means that $\omega\in Z^{2}({\mathfrak{g}};\operatorname{\mathbb{K}})$. In the case where $\omega$ is a coboundary, we call $({\mathfrak{g}},[\;.\;],\omega)$ a Frobenius Lie superalgebra, see [E, O]. Furthermore, if $\omega$ is orthosymplectic (resp. periplectic), we call $(\mathfrak{g},[\;,\;],\omega)$ an orthosymplectic (resp. periplectic) quasi-Frobenuis Lie superalgebra.
Proposition.
Let $(\mathfrak{g},[\;,\;],\omega)$ be a quasi-Frobenius Lie superalgebra and $\bullet$ be its Levi-Civita product. The following properties hold.
  1. $(i)$
    For all $u,v,w\in\mathfrak{g}$,
    (7) $\omega(\mathrm{L}_{u}^{\bullet}(v),w)=\omega(u,\mathop{\mathrm{ad}}\nolimits_{v}(w)).$
  2. $(ii)$
    We have [g, g]^= N_(g, ) = { u g, ad_u^* = -ad_u }.
  3. $(iii)$
    We have (gg)^⟂⊆[g,g]^, and Z(g) = (gg)^= N_(g, ) N_r(g, ).
In particular, $\mathfrak{g},\mathfrak{g}]^{\perp}$ is an abelian subsuperalgebra and $Z(\mathfrak{g})\subseteq[\mathfrak{g},\mathfrak{g}]^{\perp}$.
Proof.
Part (i) follows from the fact the $\omega$ is closed and Eq. (4).
Let us prove Part (ii). For all $u\in[\mathfrak{g},\mathfrak{g}]^{\perp}$ and $v,w\in\mathfrak{g}$, Part (i) implies that
$\omega(\mathrm{L}_{u}^{\bullet}(v),w)=\omega(u,[v,w])=0.$
Since $\omega$ is non-degenerate, it follows that ${\mathrm{L}}^{\bullet}_{u}=0,$ which shows that $\mathfrak{g},\mathfrak{g}]^{\perp}\subseteq N_{\ell}(\mathfrak{g},\bullet).$ The converse inclusion follows similarly. On the other hand, from Eq. (5) we know that ${\mathrm{R}}^{\bullet}_{u}$ is $\omega$-anti-symmetric; consequently, $\operatorname{ad}_{u}=-{\mathrm{R}}^{\bullet}_{u}$ is also $\omega$-anti-symmetric.
Let us prove Part (iii). Since [,]=[,][$\cdot$$\cdot$$\cdot$$\cdot$$_{\bullet}$, it is clear that $\mathfrak{g},\mathfrak{g}]\subseteq\mathfrak{g}\bullet\mathfrak{g}$ and hence $(\mathfrak{g}\bullet\mathfrak{g})^{\bot}\subseteq[\mathfrak{g},\mathfrak{g}]^{\bot}$. In addition, from Eq. (5) we obtain ${\mathrm{R}}^{\bullet}_{u}=0$, for all $u\in(\mathfrak{g}\bullet\mathfrak{g})^{\bot}$. Now, we also have $\mathrm{L}^{\bullet}_{u}=0$, for all $u\in(\mathfrak{g}\bullet\mathfrak{g})^{\perp}$, as seen in Part (i). From Eq. (5) we deduce $(\mathfrak{g}\bullet\mathfrak{g})^{\perp}\subseteq N_{\ell}(\mathfrak{g},\bullet)\cap N_{r}(\mathfrak{g},\bullet)$. Since $(\mathfrak{g}\bullet\mathfrak{g})^{\perp}=N_{r}(\mathfrak{g},\bullet)$ from Eq. (5), we finally deduce that $(\mathfrak{g}\bullet\mathfrak{g})^{\perp}=N_{\ell}(\mathfrak{g},\bullet)\cap N_{r}(\mathfrak{g},\bullet)$.
On the other hand, we have $\omega(v\bullet u,w)=\omega(v,[u,w])=0$, for all $u\in Z(\mathfrak{g})$ and $v,w\in\mathfrak{g}$. It follows that ${\mathrm{R}}^{\bullet}_{u}={\mathrm{L}}^{\bullet}_{u}=0$. Hence, $Z(\mathfrak{g})=(\mathfrak{g}\bullet\mathfrak{g})^{\perp}$. ∎
Proposition.
Let $(\mathfrak{g},[\;,\;],\omega)$ be a quasi-Frobenius nilpotent Lie superalgebra. Then $\mathfrak{g},\mathfrak{g}]$ and $Z(\mathfrak{g})$ are both degenerate.
Proof.
Since $(\mathfrak{g},[\;,\;])$ is nilpotent, then $\mathfrak{g},\mathfrak{g}]\cap Z(\mathfrak{g})\neq\{0\}$. From Prop. 3.1.3.  Proposition, we have $Z(\mathfrak{g})\subseteq[\mathfrak{g},\mathfrak{g}]^{\perp}$. It follows that $\{0\}\neq[\mathfrak{g},\mathfrak{g}]\cap Z(\mathfrak{g})\subseteq[\mathfrak{g},\mathfrak{g}]\cap[\mathfrak{g},\mathfrak{g}]^{\perp}$, and $\{0\}\neq[\mathfrak{g},\mathfrak{g}]\cap Z(\mathfrak{g})\subseteq Z(\mathfrak{g})^{\bot}\cap Z(\mathfrak{g})$, which imply that $\mathfrak{g},\mathfrak{g}]$ and $Z(\mathfrak{g})$ are degenerate. ∎
Remark.
Quasi-Frobenius Lie (super)algebras with a degenerate center have been studied in [BE, Fi].
Proposition.
Let $(\mathfrak{g},[\;,\;],\omega)$ be a quasi-Frobenius Lie superalgebra, and $\bullet$ be its Levi-Civita product. Then $(\mathfrak{g},\bullet)$ is a right symmetric superalgebra.
Proof.
Using Prop. 3.1.3.  Proposition, we have (for any $u,v,w,z\in\mathfrak{g}):$
$\omega($ $(w\bullet u)\bullet v-w\bullet(u\bullet v)-(-1)^{|u||v|}(w\bullet v)\bullet u+(-1)^{|u||v|}w\bullet(v\bullet u),z)$
$=\omega(w,-[u\bullet v-(-1)^{|u||v|}v\bullet u,z]+[u,[v,z]]-(-1)^{|u||v|}[v,[u,z]])$
$=-\omega(w,[[u,v],z]-[u,[v,z]]+(-1)^{|u||v|}[v,[u,z]])$
$=-\omega(w,[[u,v],z]+(-1)^{|u||v|+|u||z|}[[v,z],u]+(-1)^{|v||z|+|u||z|}[[z,u],v])$
$=0.$
Since $\omega$ is non-degenerate, it follows that $(\mathfrak{g},\bullet)$ is a right symmetric superalgebra. ∎
Let $(\mathfrak{g},\bullet)$ be a right-symmetric superalgebra. Consider the opposite product $\circ$ defined by
$u\circ v=-(-1)^{|u||v|}v\bullet u,$
for all $u,v\in\mathfrak{g}$. It is clear that $(\mathfrak{g},\circ)$ is a left-symmetric superalgebra. Moreover, Prop. 3.1.5.  Proposition implies that
(8) $\omega(\mathrm{L}_{u}^{\circ}(v),w)=-(-1)^{|u||v|}\omega(v,\mathop{\mathrm{ad}}\nolimits_{u}(w)),$
for all $u,v,w\in\mathfrak{g}$.
Therefore, for every quasi-Frobenius Lie superalgebra $(\mathfrak{g},[\;,\;],\omega)$, there exists a unique compatible left-symmetric structure.
Proposition.
Let $({\mathcal{A}},\circ,\omega)$ be a closed pre-symplectic superalgebra. Then the following conditions are equivalent:
  1. $(i)$
    $({\mathcal{A}},\circ)$ is a left symmetric superalgebra such that
    (9) $\omega({\mathrm{R}}_{v}^{\circ}(u),w)=(-1)^{|v||w|}\omega(u,{\mathrm{R}}_{v}^{\circ}(w)),\quad\forall u,v,w\in{\mathcal{A}}.$
  2. $(ii)$
    $({\mathcal{A}}^{-},\omega)$ is a quasi-Frobenius Lie superalgebra such that
    $\omega(\mathrm{L}_{u}^{\circ}(v),w)=-(-1)^{|u||v|}\omega(v,\mathop{\mathrm{ad}}\nolimits_{u}(w)),\quad\forall u,v,w\in{\mathcal{A}}.$
Proof.
Assume that $(i)$ holds. Since $({\mathcal{A}},\circ)$ is a left-symmetric superalgebra, it follows that $\mathcal{A}$$^{-}$ is a Lie superalgebra. Moreover, since $\omega$ is closed with respect to $\circ,$ it follows that $\omega$ is also closed with respect to $\cdot$$\cdot$$_{\circ}$. Hence, $({\mathcal{A}}^{-},\omega)$ is a quasi-Frobenius Lie superalgebra.
On the other hand, we have (for every $u,v,w\in{\mathcal{A}}$):
[…]
$=$ $-(-1)^{|u||v|}\,\omega(v,\mathop{\mathrm{ad}}\nolimits_{u}(w)).$
Conversely, assume that $(ii)$ holds. Eq. (8) implies that the product $\circ$ defines a left-symmetric structure on $\mathcal{A}$}. Moreover, for all $u,v,w\in{\mathcal{A}}$, we have
$\omega(\mathrm{L}_{u}^{\circ}(v),w)$ $=-(-1)^{|u||v|}\,\omega(v,\mathop{\mathrm{ad}}\nolimits_{u}(w))$
$=(-1)^{|u||v|+|u||w|}\,\omega(v,\mathop{\mathrm{ad}}\nolimits_{w}(u))$
$=-(-1)^{|u||v|+|u||w|+|v||w|}\,\omega(\mathrm{L}_{w}^{\circ}(v),u)$
$=(-1)^{|v||w|}\,\omega(u,\mathrm{L}_{w}^{\circ}(v)).$
This completes the proof. ∎
3.2. The symplectic product
In [BCGRS], the authors established the existence and non-uniqueness of the symplectic product in the case of Lie algebras. Here we extend the result to the super setting. The Levi-Civita product is uniquely determined by a pre-symplectic structure. In contrast, we show in Corollary 3.2.8.  Corollary that the symplectic product on a quasi-Frobenius Lie superalgebra is not unique.
Definition.
(Symplectic product) Let $(\mathfrak{g},[\;,\;],\omega)$ be a quasi-Frobenuis Lie superalgebra. A bilinear map
$\star:\mathfrak{g}\times\mathfrak{g}\to\mathfrak{g}$
is called a symplectic product if it satisfies the following conditions for all $u,v\in\mathfrak{g}$:
(10) $[u,v]_{\star}$ $=$ $[u,v],$
(11) $(\mathrm{L}_{u}^{\star})^{*}$ $=$ $-\mathrm{L}_{u}^{\star}.$
Example.
Following [BM], consider the periplectic quasi-Frobenuis Lie superalgebra $(D^{5},[\,,\,],\omega)$. In the basis: $\{e_{1},e_{2}\mid f_{1},f_{2}\}$ (even $\mid$ odd), it is defined as follows:
$e_{1},f_{1}]=f_{1},\qquad[e_{1},f_{2}]=f_{2},\qquad[e_{2},f_{2}]=f_{1},\quad\text{ and }\quad\omega=e_{1}^{*}\wedge f_{1}^{*}+e_{2}^{*}\wedge f_{2}^{*}.$
A direct computation shows that the following product is symplectic:
$\begin{array}[]{lcrlcrlcrlcrlcrlcr}e_{1}\star e_{1}&=&-\tfrac{1}{3}e_{1},&e_{1}\star e_{2}&=&-\tfrac{1}{3}e_{2},&e_{1}\star f_{1}&=&-\tfrac{1}{3}f_{1},&e_{1}\star f_{2}&=&-\tfrac{1}{3}f_{2},\\[3.0pt] e_{2}\star e_{1}&=&\tfrac{1}{3}e_{2},&e_{2}\star f_{2}&=&\tfrac{1}{3}f_{1},&f_{1}\star e_{1}&=&-\tfrac{2}{3}f_{1},&f_{2}\star e_{1}&=&-\tfrac{2}{3}f_{2},\\[3.0pt] f_{2}\star e_{2}&=&-\tfrac{2}{3}f_{2}.\end{array}$
Example.
Following [BM], consider the orthosymplectic quasi-Frobenuis Lie superalgebra $(C^{3}+A,[\,,\,],\omega)$. In the basis: $\{e_{1},e_{2}\mid f_{1},f_{2}\}$ (even $\mid$ odd), it is defined as follows:
$e_{1},f_{1}]=f_{2},\qquad[f_{1},f_{1}]=e_{1},\quad\text{ and }\quad\omega=2e_{1}^{*}\wedge e_{2}^{*}-f_{1}^{*}\wedge f_{2}^{*}.$
A direct computation shows that the following product is symplectic:
$f_{1}\star e_{1}=-f_{2},\quad f_{1}\star f_{1}=\tfrac{1}{2}e_{2}.$
We will need the following lemma later.
Lemma.
Let $({\mathcal{A}},\star,\omega)$ be a Lie-admissible superalgebra endowed with a non-degenerate bilinear anti-symmetric form such that ${\mathrm{L}}^{\star}_{u}$ is $\omega$-anti-symmetric, for any $u\in{\mathcal{A}}$. Then $({\mathcal{A}}^{-},[\,,\,]_{\star},\omega)$ is a quasi-Frobenius Lie superalgebra.
Proof.
For any $u,v,w\in{\mathcal{A}}$, we have
$\omega([u,v],w)=\omega(u\star v,w)-(-1)^{|u||v|}\omega(v\star u,w).$
By using the relation $\omega(\mathrm{L}_{u}^{\star}v,w)=-(-1)^{|u||v|}\omega(v,\mathrm{L}_{u}^{\star}w)$, one can see easily that
$\oint_{u,v,w}(-1)^{|u||w|}\omega([u,v],w)$ $=0\qed.$
Remark.
The product constructed in Lemma 3.2.4.  Lemma, whenever it exists, is a symplectic product on $({\mathcal{A}}^{-},[\,,\,]_{\star},\omega)$.
Theorem.
On a quasi-Frobenius Lie superalgebra $(\mathfrak{g},[\;,\;],\omega)$, there always exists a symplectic product.
Proof.
Suppose that $(\mathfrak{g},[\;,\;],\omega)$ is a quasi-Frobenius Lie superalgebra. Then, there exists a bilinear product $\bullet$ on $\mathfrak{g}$ compatible with the Lie bracket [,][\;,\;]; for instance, by Prop. 3.1.3.  Proposition, one can consider the Levi-Civita product associated with $(\mathfrak{g},[\;,\;],\omega)$. Define an even bilinear map $M:\mathfrak{g}\times\mathfrak{g}\to\mathfrak{g}$ by the identity:
$\omega(u\bullet v,w)-(-1)^{|u||v|}\omega(v,u\bullet w)=\omega(M(u,v),w),\quad\text{for all }u,v,w\in\mathfrak{g}.$
Since $\omega$ is anti-symmetric, we have $\omega(M(u,v),w)=(-1)^{|u||v|}\omega(v,M(u,w))$. Furthermore, the cyclicity of $\omega$ gives
$\oint_{u,v,w}(-1)^{|u||w|}\omega([u,v],w)$ $=\omega(u\bullet v,w)-(-1)^{|u||v|}\omega(v\bullet u,w)+\omega(v\bullet w,u)$
$\quad-(-1)^{|v||w|}\omega(w\bullet v,u)+\omega(w\bullet u,v)-(-1)^{|w||u|}\omega(u\bullet w,v)$
$=(-1)^{|u||w|}\omega(M(u,v),w)+(-1)^{|v||u|}\omega(M(v,w),u)$
$\quad+(-1)^{|w||v|}\omega(M(w,u),v).$
Thus, we obtain the identity $\oint_{u,v,w}(-1)^{|u||w|}\omega(M(u,v),w)=0.$
Now, define a new product $\star$ on $\mathfrak{g}$ by
$u\star v:=u\bullet v+\frac{1}{3}(M(u,v)+(-1)^{|u||v|}M(v,u)).$
We check that $\star$ is a compatible product for the Lie superalgebra $(\mathfrak{g},[\;,\;],\omega)$:
$u,v]_{\star}=[u,v]_{\bullet}+\frac{1}{3}(M(u,v)+(-1)^{|u||v|}M(v,u)-(-1)^{|u||v|}M(v,u)-M(u,v))=[u,v].$
Hence, $(\mathfrak{g},\star)$ is also compatible with the Lie superalgebra structure.
To show that $\star$ is $\omega$-anti-symmetric, we have
$\omega(u\star v,w)+(-1)^{|u||v|}\omega(v,u\star w)=\omega(u\bullet v,w)+(-1)^{|u||v|}\omega(v,u\bullet w)-\frac{1}{3}\omega(M(u,v),w)$
$\quad-\frac{1}{3}(-1)^{|u||v|}\omega(M(v,u),w))-\frac{1}{3}(-1)^{|u||v|}\omega(v,M(u,w))-\frac{1}{3}(-1)^{|u||v||u||w|}\omega(v,M(w,u)))$
$=\omega(M(u,v),w)-\frac{2}{3}\omega(M(u,v),w)-\frac{1}{3}\omega(u,M(v,w))-\frac{1}{3}(-1)^{|u||v|+|u||w|}\omega(v,M(w,u))$
$=\omega(M(u,v),w)-\frac{2}{3}\omega(M(u,v),w)-\frac{1}{3}(-1)^{|u||w|+|v||w|}\omega(w,M(u,v))$
$=(1-\frac{2}{3}-\frac{1}{3})\omega(M(u,v),w)=0.$
because $\omega(M(u,v),w)=(-1)^{|u||v|}\omega(v,M(u,w))$ and $M$ is cyclic on $\omega$. Therefore, $\star$ is $\omega$-anti-symmetric, as claimed. ∎
Lemma.
Let $E$ be a vector superspace equipped with an orthosymplectic (resp. periplectic) form $\omega$. Then there exists an even (resp. odd) symmetric bilinear map $f:E\times E\to E$ such that (for all $u,v,w\in E$)
$\omega(f(u,v),w)=-(-1)^{|u||v|}\omega(v,f(u,w)).$
Proof.
Let $(E,\omega)$ be a even (resp. odd) symplectic superspace. Consider $T$ as an even (resp. odd) symmetric trilinear form on $E$. For $u,v\in E$, define the map
$T(u,v,\cdot):E\to\mathbb{K}$
as an even (resp. odd) linear form. Since $\omega$ is non-degenerate, there exists a unique vector $f_{T}(u,v)\in E$ such that
$T(u,v,w)=\omega(f_{T}(u,v),w)\quad\text{for all }w\in E.$
Since $T$ is even (resp. odd) and symmetric on $E$, and $\omega$ is even (resp. odd), it is straightforward to show that $f_{T}:E\times E\to E$ is an even symmetric bilinear map. For any $u,v,w\in E$, we have
$\begin{array}[]{lcl}\omega(f_{T}(u,v),w)&=&T(u,v,w)=(-1)^{|v||w|}T(u,w,v)=(-1)^{|v||w|}\omega(f_{T}(u,w),v)\\[5.69054pt] &=&-(-1)^{|u||v|}\omega(v,f_{T}(u,w)).\end{array}$
This completes the proof. ∎
Proposition.
Let $(\mathfrak{g},[\;,\;],\omega)$ be a quasi-Frobenius Lie superalgebra, and let $\star$ be a symplectic product on $(\mathfrak{g},[\;,\;],\omega)$. Consider a bilinear map $f:\mathfrak{g}\times\mathfrak{g}\to\mathfrak{g}$. The product defined by
$u\bar{\star}v=u\star v+f(u,v),\quad\forall u,v\in\mathfrak{g}$
is a symplectic product on $(\mathfrak{g},[\;,\;],\omega)$ if and only if $f$ is symmetric and satisfies the identity
$\omega(f(u,v),w)=-(-1)^{|u||v|}\omega(v,f(u,w)),$
for any $u,v,w\in\mathfrak{g}.$
Proof.
Assume that $\star$ is a symplectic product on $(\mathfrak{g},[\;,\;],\omega)$. For any $u,v\in\mathfrak{g}$, we have
${}_{\bar{\star}}$ $=u\star v+f(u,v)-(-1)^{|u||v|}v\star u-(-1)^{|u||v|}f(u,v)$
$=[u,v]+f(u,v)-(-1)^{|u||v|}f(v,u).$
We therefore obtain $u,v]_{\bar{\star}}=[u,v]$ if and only if $f(u,v)=(-1)^{|u||v|}f(v,u)$.
On the other hand, for any $u,v,w\in\mathfrak{g}$, we have
$\omega(u\bar{\star}v,w)+(-1)^{|u||v|}\omega(v,u\bar{\star}w)$ $=\omega(u\star v+f(u,v),w)+(-1)^{|u||v|}\omega(v,u\star w+f(u,w))$
$=\omega(f(u,v),w)+(-1)^{|u||v|}\omega(v,f(u,w)).$
Thus, $\omega(u\bar{\star}v,w)+(-1)^{|u||v|}\omega(v,u\bar{\star}w)=0$ if and only if $\omega(f(u,v),w)=-(-1)^{|u||v|}\omega(v,f(u,w))$. ∎
Corollary.
On a quasi-Frobenius Lie superalgebra there exist many symplectic products.
Proof.
According to Prop. 3.2.5.  Theorem, there exists a symplectic product $\star$. Moreover, by Lemma 3.2.6.  Lemma, there exists an even symmetric bilinear map $f:\mathfrak{g}\times\mathfrak{g}\to\mathfrak{g}$ satisfying the condition
$\omega(f(u,v),w)=-(-1)^{|u||v|}\omega(v,f(u,w)).$
Thus, by Prop. 3.2.7.  Proposition, the product defined by
$u\bar{\star}v=u\star v+f(u,v),\quad\forall u,v\in\mathfrak{g},$
is also a symplectic product on $(\mathfrak{g},[\;,\;],\omega)$. ∎
In [BB], the authors introduced a symplectic connection that depends on the Lie bracket and the symplectic form. In this work, we introduce a natural symplectic product for quasi-Frobenius Lie superalgebras that depends on both the Lie bracket and the form, and we establish its uniqueness in Prop. 3.2.9.  Proposition.
Let $(\mathfrak{g},[\;,\;],\omega)$ be a quasi-Frobenius Lie superalgebra. Following the same procedure as in Prop. 3.2.5.  Theorem, and using Eq. (8), we define a product $\circ$ on $\mathfrak{g}$ by
$\omega(\mathrm{L}_{u}^{\circ}(v),w)=(-1)^{|u||v|}\,\omega(v,\mathop{\mathrm{ad}}\nolimits_{u}(w)),\qquad\forall u,v,w\in\mathfrak{g}.$
This product satisfies $u,v]=[u,v]_{\circ}.$ Moreover, since $\omega$ is closed, we have
$\omega(u\circ v,w)+(-1)^{|u||v|}\,\omega(v,u\circ w)=(-1)^{|u||v|}\,\omega(v\circ u,w).$
From the proof of Theorem 3.2.5.  Theorem, we consider the map $M(u,v)=-(-1)^{|u||v|}\,v\circ u.$ Using this, we define the product $\star$ on $\mathfrak{g}$ by
$u\star v=u\circ v-\tfrac{1}{3}\Big((-1)^{|u||v|}\,v\circ u+u\circ v\Big)=\tfrac{2}{3}\,u\circ v-\tfrac{1}{3}(-1)^{|u||v|}\,v\circ u.$
One can verify that this product satisfies
$\omega(u\star v,w)=\tfrac{2}{3}\,\omega(u\circ v,w)-\tfrac{1}{3}(-1)^{|u||v|}\,\omega(v\circ u,w)=-(-1)^{|u||v|}\,\tfrac{2}{3}\,\omega(v,[u,w])-\tfrac{1}{3}\,\omega(u,[v,w]).$
Since $\omega$ is closed, it follows that
(12) $\omega(u\star v,w)=\tfrac{1}{3}\,\omega([u,v],w)+\tfrac{1}{3}(-1)^{|v||w|}\,\omega([u,w],v).$
Therefore, we obtain a natural symplectic product on $\mathfrak{g}$ that depends only on the form $\omega$ and the Lie bracket.
Proposition.
Let $(\mathfrak{g},[\;,\;],\omega)$ be a quasi-Frobenius Lie superalgebra. There exists a unique natural symplectic product that depends on the product [,][\;,\;] and the form $\omega$. This product is given by (for all $u,v,w\in\mathfrak{g}$):
(13) $\omega(u\star v,w)=\frac{1}{3}\left(\omega([u,v],w)+(-1)^{|v||w|}\omega([u,w],v)\right).$
Proof.
Suppose there exists a symplectic product on $(\mathfrak{g},[\;,\;],\omega)$ such that:
$\omega(u\star v,w)=a\omega([u,v],w)+b(-1)^{|v||w|}\omega([u,w],v)+c(-1)^{|v||w|}\omega([v,w],u),$
for some scalars $a,b,c\in\mathbb{K}$. The condition $\omega(u\star v,w)=-(-1)^{|u||v|}\omega(v,u\star w)$ implies that $b=a$ and $c=0$.
Using the identity $u,v]=[u,v]_{\star}$, we have
$\omega([u,v],w)$ $=\omega(u\star v-(-1)^{|u||v|}v\star u,w)$
$=a\omega([u,v],w)+(-1)^{|v||w|}a\omega([u,w],v)-a(-1)^{|u||v|}\omega([v,u],w)$
$-(-1)^{|u||w|+|u||v|}a\omega([v,w],u).$
Since the form $\omega$ is closed, this simplifies to $\omega([u,v],w)=3a\omega([u,v],w),$ and thus $a=\frac{1}{3}$. Therefore, the unique symplectic product compatible with the bracket [,][\;,\;] and the form $\omega$ is (for all $u,v,w\in\mathfrak{g}$):
$\omega(u\star v,w)=\frac{1}{3}\left(\omega([u,v],w)-(-1)^{|v||w|}\omega([u,w],v)\right).\qed$
Definition.
We call the product defined in Eq. (13) the natural symplectic product.
4. Flat quasi-Frobenius Lie superalgebras
We investigate flat quasi-Frobenius Lie superalgebras, characterized by a natural symplectic product with identically zero curvature. This product induces a left-symmetric structure, extending the Lie algebra case studied in [BELL] to the super setting.
Definition.
Let $(\mathfrak{g},[\;,\;],\omega)$ be a quasi-Frobenius Lie superalgebra, and let $\star$ be a symplectic product on $(\mathfrak{g},[\;,\;],\omega)$. We define the curvature operator on $\mathfrak{g}$ by
$\mathcal{R}(u,v):={\mathrm{L}}^{\star}_{[u,v]}-[{\mathrm{L}}^{\star}_{u},{\mathrm{L}}^{\star}_{v}],\qquad\forall\,u,v\in\mathfrak{g},$
where ${\mathrm{L}}^{\star}_{u}$ denotes the left multiplication by $u$ with respect to $\star$.
Moreover, if this curvature operator vanishes identically, i.e.
$\mathcal{R}(u,v)=0,\quad\forall\,u,v\in\mathfrak{g},$
then $(\mathfrak{g},[\;,\;],\omega)$ is called a flat quasi-Frobenius Lie superalgebra.
It is easy to see that the curvature $\mathcal{R}$ of the symplectic product $\star$ vanishes identically if and only if $(\mathfrak{g},\star)$ is a left-symmetric superalgebra.
As we proceed in this work, we consider the natural symplectic product defined by Eq.  (13).
Let $(\mathfrak{g},[\;,\;],\omega)$ be a quasi-Frobenius Lie superalgebra, and let $\star$ be the natural symplectic product. The left and right multiplications with respect to $\star$ are expressed as (for all $u\in\mathfrak{g}$):
(14) ${\mathrm{L}}_{u}^{\star}=\tfrac{1}{3}\big(\mathop{\mathrm{ad}}\nolimits_{u}-\mathop{\mathrm{ad}}\nolimits_{u}^{*}\big),\qquad{\mathrm{R}}_{u}^{\star}=-\tfrac{1}{3}\big(2\mathop{\mathrm{ad}}\nolimits_{u}+\mathop{\mathrm{ad}}\nolimits_{u}^{*}\big),$
where $\mathop{\mathrm{ad}}\nolimits_{u}^{*}$ denotes the adjoint of $\mathop{\mathrm{ad}}\nolimits_{u}$ with respect to the form $\omega$.
Proposition.
Let $(\mathfrak{g},[\;,\;],\omega)$ be a quasi-Frobenius Lie superalgebra, and let $\star$ be the natural symplectic product. Then:
  1. (i)
    For all $u\in[\mathfrak{g},\mathfrak{g}]^{\perp}$, we have ${\mathrm{L}}_{u}^{\star}=-2{\mathrm{R}}^{\star}_{u}=\tfrac{2}{3}\,\mathop{\mathrm{ad}}\nolimits_{u}$. Moreover,
    $\mathfrak{g},\mathfrak{g}]^{\perp}=\{\,u\in\mathfrak{g}\;\mid\;({\mathrm{R}}_{u}^{\star})^{*}=-{\mathrm{R}}^{\star}_{u}\,\}.$
  2. (ii)
    The center of $\mathfrak{g}$ satisfies
    $Z(\mathfrak{g})=(\mathfrak{g}\star\mathfrak{g})^{\perp}=N_{r}(\mathfrak{g},\star)=N_{r}(\mathfrak{g},\star)\cap N_{\ell}(\mathfrak{g},\star)=N_{\ell}(\mathfrak{g},\star)\cap[\mathfrak{g},\mathfrak{g}]^{\perp}.$
Proof.
For part (i), since $\omega$ is closed, we have (for all $u,v,w\in\mathfrak{g}$):
$\omega(\mathrm{L}_{u}^{\star}v,w)=\tfrac{1}{3}\,\omega([u,v],w)+\tfrac{1}{3}(-1)^{|v||w|}\,\omega([u,w],v)=\tfrac{2}{3}\,\omega([u,v],w)+\tfrac{1}{3}(-1)^{|u||v|+|u||w|}\,\omega([v,w],u).$
Now, if $u\in[\mathfrak{g},\mathfrak{g}]^{\perp}$ and for all $v,w\in\mathfrak{g}$, we have
$\omega(\mathrm{L}_{u}^{\star}(v),w)=\tfrac{2}{3}\,\omega([u,v],w).$
Since $\omega$ is non-degenerate, it follows that ${\mathrm{L}}^{\star}_{u}=\tfrac{2}{3}\,\mathop{\mathrm{ad}}\nolimits_{u}.$ Using the relation $\mathop{\mathrm{ad}}\nolimits_{u}={\mathrm{L}}^{\star}_{u}-{\mathrm{R}}^{\star}_{u}$, we deduce ${\mathrm{R}}^{\star}_{u}=-\tfrac{1}{3}\,\mathop{\mathrm{ad}}\nolimits_{u}.$ Thus, for $u\in[\mathfrak{g},\mathfrak{g}]^{\perp}$, we obtain $({\mathrm{R}}_{u}^{\star})^{*}=-{\mathrm{R}}_{u}^{\star}$.
For part (ii), let $u\in Z(\mathfrak{g})$. Since $\mathop{\mathrm{ad}}\nolimits_{u}=0$, Eq.(14) implies that ${\mathrm{L}}^{\star}_{u}=-{\mathrm{R}}^{\star}_{u}.$ Therefore, ${\mathrm{L}}^{\star}_{u}={\mathrm{R}}^{\star}_{u}=0$.
Now, let $u\in(\mathfrak{g}\star\mathfrak{g})^{\perp}$. For all $v,w\in\mathfrak{g}$, we have
$0=\omega(v\star w,u)=-(-1)^{|v||w|}\,\omega(w,v\star u).$
Since $\omega$ is non-degenerate, this implies ${\mathrm{R}}^{\star}_{u}=0$. Moreover, $\omega([v,w],u)=0,$ for every $v,w\in\mathfrak{g},$ so that $(\mathfrak{g}\star\mathfrak{g})^{\perp}\subset[\mathfrak{g},\mathfrak{g}]^{\perp}.$
From (i), we know that for $u\in[\mathfrak{g},\mathfrak{g}]^{\perp}$, we have ${\mathrm{L}}^{\star}_{u}=-\tfrac{1}{2}{\mathrm{R}}^{\star}_{u}.$ Therefore, for $u\in(\mathfrak{g}\star\mathfrak{g})^{\perp}$, we deduce ${\mathrm{L}}^{\star}_{u}={\mathrm{R}}^{\star}_{u}=0.$ Finally, let $u\in N_{\ell}(\mathfrak{g},\star)\cap[\mathfrak{g},\mathfrak{g}]^{\perp}$. By Part (i) again, we conclude that ${\mathrm{L}}^{\star}_{u}={\mathrm{R}}^{\star}_{u}=0.$ We thus obtain the desired equalities:
$Z(\mathfrak{g})=(\mathfrak{g}\star\mathfrak{g})^{\perp}=N_{r}(\mathfrak{g},\star)\cap N_{\ell}(\mathfrak{g},\star)=N_{\ell}(\mathfrak{g},\star)\cap[\mathfrak{g},\mathfrak{g}]^{\perp}.\qed$
Let $(\mathfrak{g},[\;,\;],\omega)$ be a quasi-Frobenius Lie superalgebra. The vanishing of the curvature, $\mathcal{R}=0$, is equivalent to the relation (for all $u,v\in\mathfrak{g}$)
(15) ${\mathrm{R}}^{\star}_{u\star v}-(-1)^{|u||v|}{\mathrm{R}}^{\star}_{v}\circ{\mathrm{R}}^{\star}_{u}=[{\mathrm{L}}^{\star}_{u},{\mathrm{R}}^{\star}_{v}].$
Proposition.
Let $(\mathfrak{g},[\;,\;],\omega)$ be a flat quasi-Frobenius Lie superalgebra. For all $u,v\in[\mathfrak{g},\mathfrak{g}]^{\perp}$, we have
$\mathop{\mathrm{ad}}\nolimits_{u}\circ\mathop{\mathrm{ad}}\nolimits_{v}=0;\qquad\text{in particular,}\qquad\mathop{\mathrm{ad}}\nolimits_{u}^{2}=0.$
Proof.
Let $u,v\in[\mathfrak{g},\mathfrak{g}]^{\perp}$ and $w\in\mathfrak{g}$. From the definition of the natural symplectic product (13), we have
$\omega(u\star v,w)=\tfrac{1}{3}\,\omega([u,v],w)+\tfrac{1}{3}(-1)^{|v||w|}\,\omega([u,w],v).$
Since $v\in[\mathfrak{g},\mathfrak{g}]^{\perp}$, it follows that $\omega([u,w],v)=0$ for all $w\in\mathfrak{g}$. Hence
$\omega(u\star v,w)=\tfrac{1}{3}\,\omega([u,v],w).$
On the other hand, as $u\in[\mathfrak{g},\mathfrak{g}]^{\perp}$, we have $\mathop{\mathrm{ad}}\nolimits_{u}^{*}=-\mathop{\mathrm{ad}}\nolimits_{u}$ and therefore
$\omega([u,v],w)=-(-1)^{|u||v|}\,\omega(v,[u,w])=0,$
[…]
Using the flatness relation (see Eq. (15)), we get
$0={\mathrm{R}}^{\star}_{u\star v}-(-1)^{|u||v|}{\mathrm{R}}^{\star}_{v}\circ{\mathrm{R}}^{\star}_{u}-[{\mathrm{L}}^{\star}_{u},{\mathrm{R}}^{\star}_{v}].$
Since $u\star v=0$, and the fact that ${\mathrm{L}}^{\star}_{u}=-\tfrac{1}{2}{\mathrm{R}}^{\star}_{u}=\tfrac{2}{3}\mathop{\mathrm{ad}}\nolimits_{u}$, for $u\in[\mathfrak{g},\mathfrak{g}]^{\perp}$, by Prop. 4.2.  Proposition, this reduces to
$0=-(-1)^{|u||v|}{\mathrm{R}}^{\star}_{v}\circ{\mathrm{R}}^{\star}_{u}-[{\mathrm{L}}^{\star}_{u},{\mathrm{R}}^{\star}_{v}]=-(-1)^{|u||v|}\tfrac{1}{9}\mathop{\mathrm{ad}}\nolimits_{v}\circ\mathop{\mathrm{ad}}\nolimits_{u}+\tfrac{2}{9}[\mathop{\mathrm{ad}}\nolimits_{u},\mathop{\mathrm{ad}}\nolimits_{v}].$
Therefore,
$0=\tfrac{2}{9}\mathop{\mathrm{ad}}\nolimits_{u}\circ\mathop{\mathrm{ad}}\nolimits_{v}-\tfrac{1}{3}(-1)^{|u||v|}\mathop{\mathrm{ad}}\nolimits_{v}\circ\mathop{\mathrm{ad}}\nolimits_{u}.$
But since $u,v\in[\mathfrak{g},\mathfrak{g}]^{\perp}$, one checks that $\mathop{\mathrm{ad}}\nolimits_{u},\mathop{\mathrm{ad}}\nolimits_{v}]=\mathop{\mathrm{ad}}\nolimits_{[u,v]}=0$. Thus, the relation reduces to
$\mathop{\mathrm{ad}}\nolimits_{v}\circ\mathop{\mathrm{ad}}\nolimits_{u}=0.$
In particular, setting $u=v$, we deduce $\mathop{\mathrm{ad}}\nolimits_{u}^{2}=0.$ This completes the proof. ∎
Theorem.
Let $(\mathfrak{g},[\;,\;],\omega)$ be a flat orthosymplectic quasi-Frobenius Lie superalgebra. Then the Lie superalgebra $(\mathfrak{g},[\;,\;])$ is solvable.
Proof.
Since $\omega$ is even, the even part $(\mathfrak{g}_{\bar{0}},[\;,\;]_{\bar{0}},\omega_{\bar{0}})$ is a flat quasi-Frobenius Lie algebra. It is shown in [BELL] that such a Lie algebra is nilpotent, and hence solvable. Moreover, according to [Kac, S], the solvability of $(\mathfrak{g}_{\bar{0}},[\;,\;]_{\bar{0}})$ implies that the Lie superalgebra $(\mathfrak{g},[\;,\;])$ is solvable. Hence, $(\mathfrak{g},[\;,\;])$ is a solvable Lie superalgebra. ∎
Theorem.
Let $(\mathfrak{g},[\;,\;],\omega)$ be a flat periplectic quasi-Frobenius Lie superalgebra. Then the Lie superalgebra $(\mathfrak{g},[\;,\;])$ is nilpotent.
Proof.
Let $(\mathfrak{g}=\mathfrak{g}_{\bar{0}}\oplus\mathfrak{g}_{\bar{1}},[\;,\;],\omega)$ be a flat periplectic quasi Frobenius Lie superalgebra. According to Lemma 2.1.  Lemma, $\dim(\mathfrak{g}_{\bar{0}})=\dim(\mathfrak{g}_{\bar{1}})$. Without loss of generality, we may assume that $\mathfrak{g}$ admits a homogeneous basis (even $\mid$ odd):
$\{U_{1},\dots,U_{n}\,\mid\,V_{1},\dots,V_{n}\}.$
The Lie bracket and the periplectic form $\omega$ are completely determined by the following relations:
[…]
where all structure constants satisfy the super Jacobi identity and $\omega$ is closed.
Now, consider the Lie algebra $\bar{\mathfrak{g}}$ and the bilinear form $\bar{\omega}$ obtained by forgetting the super structure. The same set
$\{\overline{U_{1}},\dots,\overline{U_{n}},\overline{V_{1}},\dots,\overline{V_{n}}\}$
forms a basis of $\bar{\mathfrak{g}}$, and the bracket and the symplectic form are defined by
$\overline{U_{i}},\overline{U_{j}}]=\sum_{k=1}^{n}C_{ij}^{k}\overline{U_{k}},\quad[\overline{U_{i}},\overline{V_{j}}]=\sum_{k=1}^{n}E_{ij}^{k}\overline{V_{k}};\quad\bar{\omega}(\overline{U_{i}},\overline{V_{i}})=1.$
It is straightforward to check that $\bar{\mathfrak{g}}$ is a Lie algebra and $\bar{\omega}$ is closed on $\bar{\mathfrak{g}}$. In particular, $\bar{\mathfrak{g}}$ inherits a $\mathbb{Z}_{2}$-grading, so $(\bar{\mathfrak{g}},[\;,\;],\bar{\omega})$ is a quasi-Frobenius Lie algebra.
We define the natural symplectic product $\bar{\star}$ on $\bar{\mathfrak{g}}$ by
$\bar{\omega}(\bar{u}\,\bar{\star}\,\bar{v},\bar{w})=\frac{1}{3}\,\bar{\omega}([\bar{u},\bar{v}],\bar{w})+\frac{1}{3}\,\bar{\omega}([\bar{u},\bar{w}],\bar{v}),\qquad\forall\bar{u},\bar{v},\bar{w}\in\bar{\mathfrak{g}}.$
Recall that the structure constants $C_{ij}^{k}$ and $E_{ij}^{k}$ completely determine the Lie algebra structure on $\bar{\mathfrak{g}}$ and the symplectic form $\bar{\omega}$. Let us show that the product $\bar{\star}$ depends only on the Lie bracket of $\bar{\mathfrak{g}}$ and the symplectic form $\bar{\omega}$. This would imply that $(\bar{\mathfrak{g}},[\;,\;],\bar{\omega})$ is a flat quasi-Frobenius Lie algebra. Indeed,
$\omega(U_{i}\star U_{j},V_{k})=\frac{1}{3}(\omega([U_{i},U_{j}],V_{k})+\omega([U_{i},V_{k}],V_{j})$
$=\frac{1}{3}(\sum_{l=1}^{n}C_{ij}^{l}\omega(U_{l},V_{k})+\frac{1}{3}(\sum_{l=1}^{n}E_{ik}^{l}\omega(V_{l},U_{j}))=\frac{1}{3}\sum_{l=1}^{n}C_{ij}^{l}\omega(U_{l},V_{k})-\frac{1}{3}\sum_{l=1}^{n}E_{il}^{j}\omega(U_{l},V_{k})$
$=\frac{1}{3}\sum_{l=1}^{n}(C_{ij}^{l}-E_{il}^{j})\omega(U_{l},V_{k}).$
Therefore, $U_{i}\star U_{j}=\frac{1}{3}\sum_{l=1}^{n}(C_{ij}^{l}-E_{il}^{j})U_{l}$, and hence $\bar{U_{i}}\bar{\star}\bar{U_{j}}=\frac{1}{3}\sum_{l=1}^{n}(C_{ij}^{l}-E_{il}^{j})\bar{U_{l}}$. As the constant structures $C_{ij}^{k}$ and $E_{ij}^{k}$ determine completely the product $\bar{\star}$ and $\star$ is flat, it follows that $(\bar{\mathfrak{g}},[\;,\;],\bar{\omega})$ is flat.
It is shown in [BELL] that every flat quasi-Frobenius Lie algebra of this type is nilpotent. Therefore, $\bar{\mathfrak{g}}_{0}$ is nilpotent, and so is the adjoint representation of $\bar{\mathfrak{g}}_{0}$ on $\bar{\mathfrak{g}}_{1}$.
We observe that the structure constants $C_{ij}^{k}$ determine the nilpotent Lie algebra structure on $\bar{\mathfrak{g}}_{0}$, while the constants $E_{ij}^{k}$ ensure that the representation $\rho_{\bar{\mathfrak{g}}}$ is nilpotent. Consequently, $\bar{\mathfrak{g}}_{0}$ is nilpotent, and the corresponding adjoint action on $\bar{\mathfrak{g}}_{1}$ is nilpotent as well. This implies that the Lie superalgebra $\mathfrak{g}=\mathfrak{g}_{0}\oplus\mathfrak{g}_{1}$ is nilpotent [RSS]. ∎
Proposition.
Let $(\mathfrak{g},[\;,\;],\omega)$ be a flat quasi-Frobenius Lie superalgebra. Then, $\mathfrak{g},\mathfrak{g}]$ is degenerate.
Proof.
If $\omega$ is odd, then according to Theorem 4.5.  Theorem, $(\mathfrak{g},[\;,\;])$ is nilpotent. In particular, we have $\mathfrak{g},\mathfrak{g}]\cap Z(\mathfrak{g})\neq\{0\}.$ Since $\omega$ is periplectic, we also have $Z(\mathfrak{g})\subseteq[\mathfrak{g},\mathfrak{g}]^{\perp},$ and therefore
$\mathfrak{g},\mathfrak{g}]\cap[\mathfrak{g},\mathfrak{g}]^{\perp}\neq\{0\}.$
Now, suppose $\omega$ is even, and assume that $\mathfrak{g}=[\mathfrak{g},\mathfrak{g}]\oplus[\mathfrak{g},\mathfrak{g}]^{\perp}.$ According to Prop. 4.4.  Theorem, $(\mathfrak{g},[\;,\;])$ is solvable. Hence, $\mathfrak{g},\mathfrak{g}]$ is nilpotent, and its center satisfies $Z([\mathfrak{g},\mathfrak{g}])\neq\{0\}.$
We claim that $Z([\mathfrak{g},\mathfrak{g}])\subseteq Z(\mathfrak{g})$. Indeed, let $u\in Z([\mathfrak{g},\mathfrak{g}])$. For any $v\in[\mathfrak{g},\mathfrak{g}]^{\perp}$ and $w\in[\mathfrak{g},\mathfrak{g}]$, we have
$(-1)^{|u||w|}[[u,v],w]=-(-1)^{|u||v|}[[v,w],u]-(-1)^{|v||w|}[[w,u],v]=0.$
On the other hand, Prop. 4.3.  Proposition implies that $\mathrm{ad}_{v}\circ\mathrm{ad}_{w}=0$, for any $w\in[\mathfrak{g},\mathfrak{g}]^{\perp}$, and hence $[[u,v],w]=0.$ Thus $u,v]\in Z(\mathfrak{g})\subseteq[\mathfrak{g},\mathfrak{g}]^{\perp}$, which implies $[u,v]=0$. This shows that $u\in Z(\mathfrak{g})$.
Now the fact that $Z([\mathfrak{g},\mathfrak{g}])\subseteq Z(\mathfrak{g})$ and $Z(\mathfrak{g})\subseteq[\mathfrak{g},\mathfrak{g}]^{\perp}$ leads to $Z([\mathfrak{g},\mathfrak{g}])\subseteq[\mathfrak{g},\mathfrak{g}]\cap[\mathfrak{g},\mathfrak{g}]^{\perp}=\{0\}$, which is a contradiction. ∎
Proposition.
Let $(\mathfrak{g},[\;,\;],\omega)$ be a flat quasi-Frobenius Lie superalgebra. Then $N_{\ell}(\mathfrak{g},\star)$ is a two-sided ideal of $(\mathfrak{g},\star)$ and $\mathfrak{g},[\mathfrak{g},\mathfrak{g}]^{\bot}]\subseteq N_{\ell}(\mathfrak{g},\star)$.
Proof.
It is clear that $N_{\ell}(\mathfrak{g},\star)$ is a right ideal. Let us now show that it is also a left ideal. Let $u\in N_{\ell}(\mathfrak{g},\star)$ and $v\in\mathfrak{g}$. Since $(\mathfrak{g},[\;,\;],\omega)$ is flat, we have
$(-1)^{|u||v|}{\mathrm{L}}^{\star}_{v\star u}={\mathrm{L}}^{\star}_{[u,v]}=[{\mathrm{L}}^{\star}_{u},{\mathrm{L}}^{\star}_{v}]=0.$
It follows that ${\mathrm{L}}^{\star}_{v\star u}=0$, and hence $v\star u\in N_{\ell}(\mathfrak{g},\star)$.
Now, let $u,v\in\mathfrak{g}$. From (14), one has ${\mathrm{L}}^{\star}_{u}=\frac{1}{3}(\mathop{\mathrm{ad}}\nolimits_{u}-\mathop{\mathrm{ad}}\nolimits_{u}^{*}).$ Thus,
$0$ $=$ ${\mathrm{L}}^{\star}_{[u,v]}-[{\mathrm{L}}^{\star}_{u},{\mathrm{L}}^{\star}_{v}]$
$=$ $\frac{1}{3}(\mathop{\mathrm{ad}}\nolimits_{[u,v]}-\mathop{\mathrm{ad}}\nolimits_{[u,v]}^{*})-\frac{1}{9}[\mathop{\mathrm{ad}}\nolimits_{u}-\mathop{\mathrm{ad}}\nolimits_{u}^{*},\mathop{\mathrm{ad}}\nolimits_{v}-\mathop{\mathrm{ad}}\nolimits_{v}^{*}]$
$=$ $\frac{1}{3}\mathop{\mathrm{ad}}\nolimits_{[u,v]}-\frac{1}{3}\mathop{\mathrm{ad}}\nolimits_{[u,v]}^{*}-\frac{1}{9}[\mathop{\mathrm{ad}}\nolimits_{u},\mathop{\mathrm{ad}}\nolimits_{v}]+\frac{1}{9}[\mathop{\mathrm{ad}}\nolimits_{u},\mathop{\mathrm{ad}}\nolimits_{v}^{*}]+\frac{1}{9}[\mathop{\mathrm{ad}}\nolimits_{u}^{*},\mathop{\mathrm{ad}}\nolimits_{v}]-\frac{1}{9}[\mathop{\mathrm{ad}}\nolimits_{u}^{*},\mathop{\mathrm{ad}}\nolimits_{v}^{*}]$
$=$ $\frac{2}{9}\mathop{\mathrm{ad}}\nolimits_{[u,v]}-\frac{2}{9}\mathop{\mathrm{ad}}\nolimits_{[u,v]}^{*}+\frac{1}{9}[\mathop{\mathrm{ad}}\nolimits_{u}^{*},\mathop{\mathrm{ad}}\nolimits_{v}]+\frac{1}{9}[\mathop{\mathrm{ad}}\nolimits_{u},\mathop{\mathrm{ad}}\nolimits_{v}^{*}].$
Now, $\mathop{\mathrm{ad}}\nolimits_{v}^{*}=-\mathop{\mathrm{ad}}\nolimits_{v}$, for any $u\in\mathfrak{g}$ and $v\in[\mathfrak{g},\mathfrak{g}]^{\perp}$. Substituting this into the previous relation, we obtain
$0$ $=$ $\frac{2}{9}\mathop{\mathrm{ad}}\nolimits_{[u,v]}-\frac{2}{9}\mathop{\mathrm{ad}}\nolimits_{[u,v]}^{*}+\frac{1}{9}[\mathop{\mathrm{ad}}\nolimits_{u}^{*},\mathop{\mathrm{ad}}\nolimits_{v}]+\frac{1}{9}[\mathop{\mathrm{ad}}\nolimits_{u},\mathop{\mathrm{ad}}\nolimits_{v}^{*}]=\frac{1}{9}\mathop{\mathrm{ad}}\nolimits_{[u,v]}-\frac{1}{9}\mathop{\mathrm{ad}}\nolimits_{[u,v]}^{*}.$
Therefore, $\mathop{\mathrm{ad}}\nolimits_{[u,v]}=\mathop{\mathrm{ad}}\nolimits_{[u,v]}^{*}$, which implies ${\mathrm{L}}^{\star}_{[u,v]}=0$, and hence $u,v]\in N_{\ell}(\mathfrak{g},\star)$. ∎
5. Flat double extensions of flat quasi-Frobenius Lie superalgebras
In this section, we introduce the notion of flat double extension for flat quasi-Frobenius Lie superalgebras. It is inspired by the concept of double extension, originally introduced by Medina and Revoy in [MR] for quadratic Lie algebras. It is worth noticing that this notion is different from the notion of double extensions of quasi-Frobenius Lie superalgebras introduced in [BE, BM].
5.1. Central extension of left-symmetric superalgebras
Let $({\mathcal{A}},\bullet)$ be a left-symmetric superalgebra, let $\operatorname{\mathcal{H}}:=\operatorname{\mathbb{K}}e$ be a one-dimensional vector superspace, and let $\mu:{\mathcal{A}}\times{\mathcal{A}}\rightarrow\operatorname{\mathbb{K}}$ be a homogeneous bilinear map of a given parity. We define a new product $\diamond$ on the vector space $\widetilde{{\mathcal{A}}}:={\mathcal{A}}\oplus\operatorname{\mathbb{K}}e$ as follows:
$(u+\alpha e)\diamond(v+\beta e):=u\bullet v+\mu(u,v)e,\qquad\forall\,u,v\in{\mathcal{A}},\;\alpha,\beta\in\operatorname{\mathbb{K}}.$
[…]
(16) $\mu(\mathrm{L}_{u}^{\bullet}(v),w)-\mu(u,\mathrm{L}_{v}^{\bullet}(w))=(-1)^{|u||v|}\big(\mu(\mathrm{L}_{v}^{\bullet}(u),w)-\mu(v,\mathrm{L}_{u}^{\bullet}(w))\big),$
[…]
5.2. Almost semi-direct product of left-symmetric superalgebras
Let $({\mathcal{A}},\bullet)$ be a left-symmetric superalgebra. Let $\mathcal{S}:=\operatorname{\mathbb{K}}d$ be a one-dimensional vector superspace, let $D:{\mathcal{A}}\rightarrow{\mathcal{A}},\;\xi:{\mathcal{A}}\rightarrow{\mathcal{A}},$ be homogeneous endomorphisms of parity $|d|$, and let $b_{0}\in{\mathcal{A}}_{\bar{0}}$ and a scalar $\lambda\in\operatorname{\mathbb{K}}$. On the vector superspace $\overline{{\mathcal{A}}}:=\operatorname{\mathbb{K}}d\oplus{\mathcal{A}},$ we define a new product $\bar{\bullet}$ as follows: (here $\lambda=0$ if $d$ is odd)
$d\bar{\bullet}d:=\lambda d+b_{0},\qquad d\bar{\bullet}u:=D(u),\qquad u\bar{\bullet}d:=\xi(u),\qquad u\bar{\bullet}v:=u\bullet v,$
for all $u,v\in{\mathcal{A}}$.
Then $(\overline{{\mathcal{A}}},\bar{\bullet})$ is a left-symmetric superalgebra if and only if the following conditions are satisfied (for all $u,v\in{\mathcal{A}}$):
(17) $\begin{cases}\xi([u,v]_{\bullet})=\mathrm{L}_{u}^{\bullet}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\bullet}(\xi(u)),\\[3.99994pt] D(\mathrm{L}_{u}^{\bullet}(v))=\mathrm{L}_{D(u)}^{\bullet}(v)+(-1)^{|d||u|}\mathrm{L}_{u}^{\bullet}(D(v))-(-1)^{|d||u|}\mathrm{L}_{\xi(u)}^{\bullet}(v),\\[3.99994pt] (\xi\circ D-D\circ\xi)(u)=(-1)^{|d||u|}\big(\xi^{2}-\lambda\xi-{\mathrm{R}}^{\bullet}_{b_{0}}\big)(u),\\[3.99994pt] (1-(-1)^{|d|})\big(D^{2}-\lambda D-{\mathrm{L}}^{\bullet}_{b_{0}}\big)=0,\;(1-(-1)^{|d|})(\xi-D)(b_{0})=0.\end{cases}$
Definition.
If $(D,\xi,b_{0},\lambda)$ satisfies the compatibility conditions (17), then the left-symmetric superalgebra $(\overline{{\mathcal{A}}},\bar{\bullet})$ is called the almost semi-direct product of the left-symmetric superalgebra $({\mathcal{A}},\bullet)$ by the one-dimensional vector superspace $\operatorname{\mathbb{K}}d$ by means of $(D,\xi,b_{0},\lambda)$. In this case, the quadruple $(\mathcal{S},{\mathcal{A}},D,\xi,b_{0},\lambda)$ is referred to as a context of almost semi-direct product of left-symmetric superalgebras.
Remark.
[…]
5.5. Flat double extensions of flat quasi-Frobenius Lie superalgebras
In this paragraph, we introduce the notion of a flat double extension of flat quasi-Frobenius Lie superalgebras. We then show that if a flat orthosymplectic quasi-Frobenius Lie superalgebra is nilpotent, then any its flat double extension of is also nilpotent.
5.5.1. Even flat double extensions of orthosymplectic quasi-Frobenius Lie superalgebras.
Let $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ be a flat orthosymplectic quasi-Frobenius Lie superalgebra. Denote by $\star$ the natural symplectic product associated with it. Let $V=\operatorname{\mathbb{K}}d$ be a one-dimensional vector superspace and $V^{*}=\operatorname{\mathbb{K}}e$ be its dual.
In order to introduce the process of flat double extension by $V$, we consider an even linear map $\xi:\mathfrak{b}\rightarrow\mathfrak{b},$ and an even element $b_{0}\in\mathfrak{b}_{\bar{0}}$. We extend the linear maps $\xi$ and $D:=\xi^{*}-\xi$ to $\widetilde{D},\widetilde{\xi}:\mathfrak{b}\oplus V^{*}\rightarrow\mathfrak{b}\oplus V^{*},$ as follows:
(18) $\begin{array}[]{c}\widetilde{D}(u+\alpha e):=(\xi^{*}-\xi)(u)+\dfrac{1}{3}\,\omega(b_{0},u)\,e,\quad\widetilde{\xi}(u+\alpha e):=\xi(u)-\dfrac{2}{3}\,\omega(b_{0},u)\,e.\end{array}$
for all $u\in\mathfrak{b}$ and $\alpha\in\operatorname{\mathbb{K}}$. We also choose $\widetilde{b_{0}}:=\dfrac{1}{3}\,b_{0}\in\mathfrak{b}\oplus V^{*}$.
Lemma.
The superspace $\mathcal{H}:=\mathfrak{b}\oplus V^{*}$ equipped with the product
$(u+\alpha e)\diamond(v+\beta e):=u\star v+\omega(\xi(u),v)\,e,\qquad\forall\,u,v\in\mathfrak{b},\;\alpha,\beta\in\operatorname{\mathbb{K}},$
is a left-symmetric superalgebra if and only if
(19) $\xi([u,v]_{\mathfrak{b}})=\mathrm{L}_{u}^{\star}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\star}(\xi(u)),$
for every $u,v\in\mathfrak{b}$.
Proof.
The superalgebra $(\mathcal{H},\diamond)$ is left-symmetric if and only if the bilinear map
$\mu(u,v):=\omega(\xi(u),v)$
satisfies Eq. (16), which is equivalent to Eq. (19). ∎
Now, let us also assume that $(V,\mathfrak{b},\xi^{*}-\xi,\xi^{*},\tfrac{1}{3}b_{0},0)$ is a context of almost semi-direct product of left-symmetric superalgebras.
Lemma.
$(V,\mathcal{H}:=\mathfrak{b}\oplus V^{*},\widetilde{D},\widetilde{\xi},\widetilde{b_{0}},0)$ is a context of almost semi-direct product of left-symmetric superalgebras if and only if for all $u,v\in\mathfrak{b}$, the following conditions hold:
(20) $\left\{\begin{array}[]{lcllcl}\xi([u,v])&=&\mathrm{L}_{u}^{\star}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\star}(\xi(u)),&\xi^{*}\circ\xi&=&\tfrac{1}{3}({\mathrm{R}}^{\star}_{b_{0}}+({\mathrm{R}}^{\star}_{b_{0}})^{*}),\\[5.69054pt] D(u\star v)&=&D(u)\star v+u\star D(v)-\xi(u)\star v,&[\xi,\xi^{*}]&=&\xi^{2}-\tfrac{1}{3}{\mathrm{R}}^{\star}_{b_{0}},\;b_{0}\in\ker{D},\end{array}\right$.
where $D=\xi^{*}-\xi$.
Proof.
Let $u,v\in\mathfrak{b}$ and $\alpha,\beta\in\mathbb{K}$. A straightforward computation shows that the equality
$\widetilde{\xi}([u,v])=\mathrm{L}_{u}^{\diamond}(\widetilde{\xi}(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\diamond}(\widetilde{\xi}(u))$
is equivalent to
$\xi([u,v])=\mathrm{L}_{u}^{\star}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\star}(\xi(u)),\quad\text{and}\quad\xi^{*}\circ\xi=\tfrac{1}{3}({\mathrm{R}}^{\star}_{b_{0}}+({\mathrm{R}}^{\star}_{b_{0}})^{*}).$
Moreover, the relation
$\widetilde{D}(u)\diamond v-\widetilde{D}(u\diamond v)=(-1)^{|d||u|}\big(\widetilde{\xi}(u)\diamond v-u\diamond\widetilde{D}(v)\big)$
is equivalent to
$D(u\star v)=D(u)\star v+u\star D(v)-\xi(u)\star v,$
together with
$\xi^{*},\xi]=\xi^{2}-\tfrac{1}{3}{\mathrm{R}}_{b_{0}}^{\star}.$
Finally, the identity
$\widetilde{\xi},\widetilde{D}](u)=(-1)^{|d||u|}\big(\widetilde{\xi}^{2}-{\mathrm{R}}^{\diamond}_{\widetilde{b_{0}}}\big)(u)$
is equivalent to
$\xi,\xi^{*}]=\xi^{2}-\tfrac{1}{3}{\mathrm{R}}_{b_{0}}^{\star},\quad\text{and}\quad b_{0}\in\ker(\xi^{*}-\xi).$
Since $(V,\mathfrak{b},\xi^{*}-\xi,\xi,\tfrac{1}{3}b_{0},0)$ is a context of almost semi-direct product of left-symmetric superalgebras, it follows that $(V,\mathcal{H}:=\mathfrak{b}\oplus V^{*},\widetilde{D},\widetilde{\xi},\widetilde{b_{0}},0)$ is also a context of almost semi-direct product of left-symmetric superalgebras if and only if the conditions (20) are satisfied. This proves the lemma. ∎
Remark.
Let $(V,\mathcal{H}:=\mathfrak{b}\oplus V^{*},\widetilde{D},\widetilde{\xi},\widetilde{b_{0}},0)$ be a context of an almost semi-direct product of left-symmetric superalgebras. The endomorphism defined by $\Delta(u):=\xi^{*}-2\xi$, for all $u\in\mathfrak{b}$, is a derivation on $(\mathfrak{b},[\,,]_{\mathfrak{b}})$.
We now define the even flat double extension of a flat orthosymplectic quasi-Frobenius Lie superalgebra by the one-dimensional Lie algebra.
Theorem.
Let $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ be a flat orthosymplectic quasi-Frobenius Lie superalgebra, and let $\star_{\mathfrak{b}}$ be the natural symplectic product associated with it. Let $V=\mathbb{K}d$ be a one-dimensional vector space and $V^{*}=\mathbb{K}e$ be its dual. Assume that there exist an even linear map $\xi:\mathfrak{b}\to\mathfrak{b}$, and an element $b_{0}\in\mathfrak{b}_{\bar{0}}$, satisfying the system (20). Define $\mathfrak{g}:=\mathbb{K}d\oplus\mathfrak{b}\oplus\mathbb{K}e,$ equipped with the Lie bracket and even bilinear form $\omega$ as follows:
(21) $d,u]=(\xi^{*}-2\xi)(u)+\omega_{\mathfrak{b}}(b_{0},u)e,\quad\text{ and }\quad[u,v]=[u,v]_{\mathfrak{b}}+\omega_{\mathfrak{b}}\big((\xi+\xi^{*})(u),v\big)e,$
for all $u,v\in\mathfrak{b}$ and
$\omega|_{\mathfrak{b}\times\mathfrak{b}}=\omega_{\mathfrak{b}},\qquad\omega(e,d)=-\omega(d,e)=1,\qquad\omega(d,\mathfrak{b})=\omega(e,\mathfrak{b})=0.$
Then, $(\mathfrak{g},[\;,\;],\omega)$ is a flat orthosymplectic quasi-Frobenius Lie superalgebra.
Furthermore, the natural symplectic product $\star$ associated with $(\mathfrak{g},[\;,\;],\omega)$, is given by
(22) $\begin{array}[]{lll}\displaystyle e\star u=u\star e=0,&e\star d=d\star e=0,&\displaystyle d\star u=(\xi^{*}-\xi)(u)+\tfrac{1}{3}\omega_{\mathfrak{b}}(b_{0},u)e,\\[3.00003pt] \displaystyle u\star d=\xi(u)-\tfrac{2}{3}\omega_{\mathfrak{b}}(b_{0},u)e,&d\star d=\tfrac{1}{3}b_{0},&u\star v=u\star_{\mathfrak{b}}v+\omega_{\mathfrak{b}}(\xi(u),v)e,\end{array}$
for all $u,v\in\mathfrak{b}$.
The flat orthosymplectic quasi-Frobenius Lie superalgebra $(\mathfrak{g},[\;,\;],\omega)$ is called the even flat double extension of $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ by means of $(\xi,b_{0}).$
Proof.
It is easy to check that $(\mathfrak{g},[\;,\;],\omega)$ is an orthosymplectic quasi-Frobenius Lie superalgebra. Now, we show that it is flat. From Eq. (13), we derive the natural symplectic product given by (22). Hence
$(u+\alpha e)\star(v+\beta e)=u\star_{\mathfrak{b}}v+\omega_{\mathfrak{b}}(\xi(u),v)e,$
for all elements $u,v\in\mathfrak{b}$ and $\alpha,\beta\in\mathbb{K}$. By Lemma 5.5.2.  Lemma, the first condition of system (20), for all $u,v\in\mathfrak{b}$
$\xi([u,v])=\mathrm{L}_{u}^{\star_{\mathfrak{b}}}(v)-(-1)^{|u||v|}\mathrm{L}_{v}^{\star_{\mathfrak{b}}}(u),$
implies that the above product defines on the vector superspace $\mathcal{H}=\mathfrak{b}\oplus\mathbb{K}e$ a left-symmetric structure which is a central extension of $\mathfrak{b}$ by $V$ by means of $\mu$, where $\mu(u,v)=\omega_{\mathfrak{b}}(\xi(u),v)$, for all $u,v\in\mathfrak{b}$.
Now, since $(\xi,b_{0})$ satisfies the system (20), it follows from Lemma 5.5.3.  Lemma, that $(V,\mathcal{H}:=\mathfrak{b}\oplus V^{*},\widetilde{D},\widetilde{\xi},\widetilde{b_{0}},0)$ is a context of an almost semi-direct product of left-symmetric superalgebras. Hence, $(\mathfrak{g},\star)$ is a left-symmetric superalgebra, which implies that $(\mathfrak{g},[\,,\,],\omega)$ is a flat orthosymplectic quasi-Frobenius Lie superalgebra. ∎
Lemma.
Let $(\mathfrak{g},[\;,\;],\omega)$ be an orthosymplectic quasi-Frobenius nilpotent Lie superalgebra. If $a\in\mathfrak{g}_{\bar{0}}$, then $\mathop{\mathrm{tr}}\nolimits({\mathrm{R}}^{\star}_{a})=0$.
Proof.
Let $a\in\mathfrak{g}_{\bar{0}}$. Then, we have $\omega({\mathrm{L}}^{\star}_{a}(u),v)=-\,\omega(u,{\mathrm{L}}^{\star}_{a}(v)),$ for all $u,v\in\mathfrak{g}$. Choose a homogeneous basis (even $\mid$ odd) $\{X_{1},Y_{1},\ldots,X_{n},Y_{n},\mid f_{1},\ldots,f_{m}\}$ of $\mathfrak{g}$ such that
$\omega(X_{i},Y_{i})=-\omega(Y_{i},X_{i})=-1,\quad\text{and}\quad\omega(f_{i},f_{i})=\epsilon_{i},\ \text{where }\epsilon_{i}=\pm 1.$
We compute the (ordinary) trace of ${\mathrm{L}}^{\star}_{a}$:
$\mathop{\mathrm{tr}}\nolimits({\mathrm{L}}^{\star}_{a})$ $=\sum_{i=1}^{n}\omega({\mathrm{L}}^{\star}_{a}(X_{i}),Y_{i})-\sum_{i=1}^{n}\omega({\mathrm{L}}^{\star}_{a}(Y_{i}),X_{i})+\sum_{i=1}^{m}\epsilon_{i}\,\omega({\mathrm{L}}^{\star}_{a}(f_{i}),f_{i})$
$=-\sum_{i=1}^{n}\omega(X_{i},{\mathrm{L}}^{\star}_{a}(Y_{i}))+\sum_{i=1}^{n}\omega(Y_{i},{\mathrm{L}}^{\star}_{a}(X_{i}))-\sum_{i=1}^{m}\epsilon_{i}\,\omega(f_{i},{\mathrm{L}}^{\star}_{a}(f_{i}))$
$=\sum_{i=1}^{n}\omega({\mathrm{L}}^{\star}_{a}(Y_{i}),X_{i})-\sum_{i=1}^{n}\omega({\mathrm{L}}^{\star}_{a}(X_{i}),Y_{i})-\sum_{i=1}^{m}\epsilon_{i}\,\omega({\mathrm{L}}^{\star}_{a}(f_{i}),f_{i})$
$=-\,\mathop{\mathrm{tr}}\nolimits({\mathrm{L}}^{\star}_{a}).$
Hence $\mathop{\mathrm{tr}}\nolimits({\mathrm{L}}^{\star}_{a})=0$. Since $\mathop{\mathrm{ad}}\nolimits_{a}={\mathrm{L}}^{\star}_{a}-{\mathrm{R}}^{\star}_{a}$, we have
$\mathop{\mathrm{tr}}\nolimits(\mathop{\mathrm{ad}}\nolimits_{a})=\mathop{\mathrm{tr}}\nolimits({\mathrm{L}}^{\star}_{a})-\mathop{\mathrm{tr}}\nolimits({\mathrm{R}}^{\star}_{a})=-\mathop{\mathrm{tr}}\nolimits({\mathrm{R}}^{\star}_{a}).$
But $(\mathfrak{g},[\;,\;])$ is nilpotent, so $\mathop{\mathrm{tr}}\nolimits(\mathop{\mathrm{ad}}\nolimits_{a})=0$. Therefore $\mathop{\mathrm{tr}}\nolimits({\mathrm{R}}^{\star}_{a})=0$. ∎
Lemma.
Let $(\mathfrak{g},[\;,\;],\omega)$ be a flat orthosymplectic quasi-Frobenius nilpotent Lie superalgebra. If $a\in\mathfrak{g}_{\bar{0}}$, then the right multiplication operator ${\mathrm{R}}_{a}$ is nilpotent.
Proof.
Since $(\mathfrak{g},[\;,\;],\omega)$ is a flat, then according to relation (14), we have
(23) ${\mathrm{R}}^{\star}_{u\star v}-(-1)^{|u||v|}{\mathrm{R}}^{\star}_{v}\circ{\mathrm{R}}^{\star}_{u}=[{\mathrm{L}}^{\star}_{u},{\mathrm{R}}^{\star}_{v}].$
Now, let $a\in\mathfrak{g}_{0}$. We define inductively $x_{0}=a,\;x_{n+1}={\mathrm{L}}^{\star}_{x_{n}}(a),\quad\text{for any }n\geq 0.$ We shall show by induction that, for any $n\geq 1$, we have
(24) ${\mathrm{R}}^{\star}_{x_{n}}=({\mathrm{R}}^{\star}_{a})^{\,n+1}+({\mathrm{R}}^{\star}_{a})^{\,n-2}\circ[{\mathrm{L}}^{\star}_{a},{\mathrm{R}}^{\star}_{a}]+({\mathrm{R}}^{\star}_{a})^{\,n-3}\circ[{\mathrm{L}}^{\star}_{x_{1}},{\mathrm{R}}^{\star}_{a}]+\dots+[{\mathrm{L}}^{\star}_{x_{n-1}},{\mathrm{R}}^{\star}_{a}].$
The relation is clearly true for $n=1$. Assume it holds for some $n\geq 1$. Then, using Eq. (23), we have
${\mathrm{R}}^{\star}_{x_{n+1}}$ $={\mathrm{R}}^{\star}_{x_{n}\star a}={\mathrm{R}}^{\star}_{a}\circ{\mathrm{R}}^{\star}_{x_{n}}+[{\mathrm{L}}^{\star}_{x_{n}},{\mathrm{R}}^{\star}_{a}]$
$={\mathrm{R}}^{\star}_{a}\circ\Big(({\mathrm{R}}^{\star}_{a})^{\,n+1}+({\mathrm{R}}^{\star}_{a})^{\,n-2}\circ[{\mathrm{L}}^{\star}_{a},{\mathrm{R}}^{\star}_{a}]+({\mathrm{R}}^{\star}_{a})^{\,n-3}\circ[{\mathrm{L}}^{\star}_{x_{1}},{\mathrm{R}}^{\star}_{a}]+\dots+[{\mathrm{L}}^{\star}_{x_{n-1}},{\mathrm{R}}^{\star}_{a}]\Big)+[{\mathrm{L}}^{\star}_{x_{n}},{\mathrm{R}}^{\star}_{a}]$
$=({\mathrm{R}}^{\star}_{a})^{\,n+2}+({\mathrm{R}}^{\star}_{a})^{\,n-1}\circ[{\mathrm{L}}^{\star}_{a},{\mathrm{R}}^{\star}_{a}]+({\mathrm{R}}^{\star}_{a})^{\,n-2}\circ[{\mathrm{L}}^{\star}_{x_{1}},{\mathrm{R}}^{\star}_{a}]+\dots+[{\mathrm{L}}^{\star}_{x_{n}},{\mathrm{R}}^{\star}_{a}].$
This completes the induction step.
On the other hand, since $(\mathfrak{g},[\;,\;])$ is nilpotent, Lemma 5.5.6.  Lemma implies that $\mathop{\mathrm{tr}}\nolimits({\mathrm{R}}^{\star}_{b})=0,$ for all $b\in\mathfrak{g}_{0}$. Now, we put $b=x_{n+1}\in\mathfrak{g}_{0}$, we get
$\operatorname{tr}({\mathrm{R}}^{\star}_{x_{n+1}})=0\quad\text{and}\quad\operatorname{tr}(({\mathrm{R}}^{\star}_{a})^{k}\circ[{\mathrm{L}}^{\star}_{a},{\mathrm{R}}^{\star}_{a}])=0\quad\text{for all }n\,,k\in\mathbb{N},$
since,
$\operatorname{tr}(({\mathrm{R}}^{\star}_{a})^{k}\circ[{\mathrm{L}}^{\star}_{a},{\mathrm{R}}^{\star}_{a}])$ $=\operatorname{tr}(({\mathrm{R}}^{\star}_{a})^{k}\circ({\mathrm{L}}^{\star}_{a}\circ{\mathrm{R}}^{\star}_{a}-{\mathrm{R}}^{\star}_{a}\circ{\mathrm{L}}^{\star}_{a})=\operatorname{tr}(({\mathrm{R}}^{\star}_{a})^{k}\circ({\mathrm{L}}^{\star}_{a}\circ{\mathrm{R}}^{\star}_{a}))-\operatorname{tr}(({\mathrm{R}}^{\star}_{a})^{k+1}\circ{\mathrm{L}}^{\star}_{a}))$
$=\operatorname{tr}(({\mathrm{R}}^{\star}_{a})^{k+1}\circ{\mathrm{L}}^{\star}_{a})-\operatorname{tr}(({\mathrm{R}}^{\star}_{a})^{k+1}\circ{\mathrm{L}}^{\star}_{a})=0.$
From Eq. (24), we deduce by induction that
$\operatorname{tr}(({\mathrm{R}}^{\star}_{a})^{n+1})=0\quad\text{for all }n\in\mathbb{N},$
and hence ${\mathrm{R}}^{\star}_{a}$ is nilpotent. ∎
Theorem.
Let $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ be a flat orthosymplectic quasi-Frobenius Lie superalgebra and let $(\mathfrak{g},[\;,\;],\omega)$ be the even flat double extension of $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ by means of $(\xi,b_{0})$. If $(\mathfrak{b},[\;,\;]_{\mathfrak{b}})$ is nilpotent, then $(\mathfrak{g},[\;,\;])$ is nilpotent as well.
Proof.
First, we show that, for any $k\in\mathbb{N}^{*}$ and $a\in\mathfrak{b}_{0}$,
[…]
Indeed, for any $a\in\mathfrak{b}_{\bar{0}}$ and $u\in\mathfrak{b}$, according to system (20), we have
$\xi([u,a])=u\star\xi(a)-a\star\xi(u),$
which implies
$\xi\circ{\mathrm{R}}^{\star_{\mathfrak{b}}}_{a}-\xi\circ{\mathrm{L}}^{\star_{\mathfrak{b}}}_{a}={\mathrm{R}}^{\star_{\mathfrak{b}}}_{\xi(a)}-{\mathrm{L}}^{\star_{\mathfrak{b}}}_{a}\circ\xi.$
[…]
$\xi^{k+1}([a,u])=\xi^{k+1}\circ{\mathrm{L}}^{\star_{\mathfrak{b}}}_{a}(u)-\xi^{k+1}\circ{\mathrm{R}}^{\star_{\mathfrak{b}}}_{a}(u).$
On the other hand, for any $k\in\mathbb{N}$, we have
$\xi^{k+1}([a,u])=\xi^{k}\circ\xi([a,u])=\xi^{k}\big(a\star\xi(u)-u\star\xi(a)\big)=\xi^{k}\circ{\mathrm{L}}^{\star_{\mathfrak{b}}}_{a}\circ\xi(u)-\xi^{k}\circ{\mathrm{R}}^{\star_{\mathfrak{b}}}_{\xi(a)}(u).$
Comparing the two expressions and taking the trace, we obtain
[…]
which proves the property for $k+1$.
We set $\Delta:=\xi^{*}-2\xi$. From (20), we have $\xi,\Delta]=\xi^{2}-\frac{1}{3}{\mathrm{R}}^{\star_{\mathfrak{b}}}_{b_{0}}.$ By induction, one can show that for any $k\in\mathbb{N}^{*}$, we have
$\xi^{k},\Delta]=k\xi^{k+1}-\frac{1}{3}\sum_{p=0}^{k-1}\xi^{p}\circ{\mathrm{R}}^{\star_{\mathfrak{b}}}_{b_{0}}\circ\xi^{k-1-p}.$
[…]
From the system (20), we have $\xi^{*}\circ\xi=\frac{1}{3}({\mathrm{R}}^{\star_{\mathfrak{b}}}_{b_{0}}+({\mathrm{R}}^{\star_{\mathfrak{b}}}_{b_{0}})^{*}),$ and $\xi,\xi^{*}]=\xi^{2}-\frac{1}{3}{\mathrm{R}}^{\star_{\mathfrak{b}}}_{b_{0}}.$ It follows that $\xi\circ\xi^{*}=\xi^{2}+\frac{1}{3}({\mathrm{R}}^{\star_{\mathfrak{b}}}_{b_{0}})^{*},$ and $(\xi^{*})^{2}=[\xi,\xi^{*}]+\frac{1}{3}({\mathrm{R}}_{b_{0}}^{\star_{\mathfrak{b}}})^{*}$.
Replacing these into the expression $\Delta^{2}=(\xi^{*}-2\xi)^{2},$ we get
$\Delta^{2}=3\xi^{2}-3\xi^{*}\circ\xi=-3\Delta\circ\xi-3\xi^{2}.$
By induction, one can show that $\Delta^{k}=a_{k}\Delta\circ\xi^{k-1}+b_{k}\xi^{k}$ for all $k\geq 2$, where $a_{k},b_{k}$ are some scalars. Since $\xi$ is nilpotent, it follows that $\Delta$ is nilpotent.
Now, if $(\mathfrak{g},[\;,\;],\omega)$ is an even flat double extension of $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$, the Lie bracket in $\mathfrak{g}$ is given by
$d,u]=\Delta(u)+\omega_{\mathfrak{b}}(b_{0},u)e,\qquad[u,v]=[u,v]_{\mathfrak{b}}+\omega_{\mathfrak{b}}\big((\xi+\xi^{*})(u),v\big)e,$
for all $u,v\in\mathfrak{b}$. By induction, one can show that for any $k\in\mathbb{N}$, we have
$\mathop{\mathrm{ad}}\nolimits^{k}_{d}(u)=\Delta^{k}(u)+\omega_{\mathfrak{b}}(b_{0},\Delta^{k-1}(u))e,\text{ and }\mathop{\mathrm{ad}}\nolimits_{u}^{k}(v)=(\mathop{\mathrm{ad}}\nolimits_{u}^{\mathfrak{b}})^{k}(v)+\omega_{\mathfrak{b}}((\xi+\xi^{*})(u),(\mathop{\mathrm{ad}}\nolimits_{u}^{\mathfrak{b}})^{k-1}(v))e,$
for all $u,v\in\mathfrak{g}$. Since $(\mathfrak{b},[\;,\;]_{\mathfrak{b}})$ is nilpotent and $\Delta$ is nilpotent, it follows that $(\mathfrak{g},[\;,\;])$ is nilpotent. This completes the proof. ∎
5.5.9. Odd flat double extensions of orthosympletic Lie superalgebras
We now introduce the notion of odd flat double extension of flat orthosymplectic quasi-Frobenius Lie superalgebras by a purely odd one-dimensional vector superspace.
Let $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ be a flat orthosymplectic quasi-Frobenius Lie superalgebra, and denote by $\star$ the natural symplectic product associated with it. Let $V=\operatorname{\mathbb{K}}d$ be a purely odd one-dimensional vector superspace and $V^{*}=\operatorname{\mathbb{K}}e$ be its dual.
In order to introduce the process of the odd flat double extension, we consider an odd linear map $\xi:\mathfrak{b}\rightarrow\mathfrak{b},$ and an even element $b_{0}\in\mathfrak{b}_{\bar{0}}$. For all $u\in{\mathfrak{b}}$, we define the linear map
(26) $D(u)=-(-1)^{|u|}(\xi^{*}+\xi)(u).$
We extend the two maps $D$ and $\xi$ to $\widetilde{D},\widetilde{\xi}:\mathfrak{b}\oplus V^{*}\longrightarrow\mathfrak{b}\oplus V^{*},$ as follows:
(27) $\widetilde{D}(u+\alpha e)=-(-1)^{|u|}(\xi^{*}+\xi)(u)+\,\omega(b_{0},u)\,e,\quad\text{ and }\quad\widetilde{\xi}(u+\alpha e)=\xi(u),$
for all $u\in\mathfrak{b}$ and $\alpha\in\operatorname{\mathbb{K}}$. Put $\widetilde{b_{0}}=b_{0}.$
Lemma.
The superspace $\mathcal{H}:=\mathfrak{b}\oplus V^{*}$ equipped with the product
$(u+\alpha e)\diamond(v+\beta e):=u\star v+(-1)^{|v|}\omega(\xi(u),v)\,e,\qquad\forall\,u,v\in\mathfrak{b},\;\alpha,\beta\in\operatorname{\mathbb{K}},$
is a left-symmetric superalgebra if and only if
(28) $\xi([u,v])=\mathrm{L}_{u}^{\star}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\star}(\xi(u)),$
for all $u,v\in\mathfrak{b}$.
Proof.
The algebra $(\mathcal{H},\diamond)$ is a left-symmetric superalgebra if and only if the bilinear map
$\mu(u,v):=\omega(\xi(u),v)$
satisfies condition (16), which is equivalent to relation (28).∎
Now, let us also assume that $(V,\mathfrak{b},D,\xi,b_{0},0)$ is a context of almost semi-direct product of left-symmetric superalgebras, where $D(u)=-(-1)^{|u|}(\xi^{*}+\xi)(u)$, for all $u\in\mathfrak{b}$.
Lemma.
$(V,\mathcal{H}:=\mathfrak{b}\oplus V^{*},\widetilde{D},\widetilde{\xi},\widetilde{b_{0}},0)$ is a context of almost semi-direct product of left-symmetric superalgebras if and only if
(29) $\left\{\begin{array}[]{rclrcl}\xi([u,v])&=&\mathrm{L}_{u}^{\star}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\star}(\xi(u)),&\left[\xi,\xi^{*}\right]&=&{\mathrm{R}}^{\star}_{b_{0}}-3\xi^{2},\\[5.69054pt] D(u\star v)&=&D(u)\star v+(-1)^{|u|}u\star D(v)+\xi(u)\star v,&{\mathrm{L}}^{\star}_{b_{0}}&=&-(\xi+\xi^{*})^{2},\\[5.69054pt] (2\xi+\xi^{*})(b_{0})&=&0,\end{array}\right$.
for all $u,v\in\mathfrak{b}$.
Proof.
Let $u,v\in\mathfrak{b}$ and $\alpha,\beta\in\mathbb{K}$. A straightforward computation shows that the equality
$\widetilde{\xi}([u,v])=\mathrm{L}_{u}^{\diamond}(\widetilde{\xi}(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\diamond}(\widetilde{\xi}(u))$
is equivalent to
$\xi([u,v])=\mathrm{L}_{u}^{\star}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\star}(\xi(u)).$
Moreover, the relation
$\widetilde{D}(u)\diamond v-\widetilde{D}(u\diamond v)=(-1)^{|d||u|}\big(\widetilde{\xi}(u)\diamond v-u\diamond\widetilde{D}(v)\big)$
is equivalent to
$(-1)^{|u|}(\xi^{*}+\xi)(u)\star v+(-1)^{|u|+|v|}(\xi^{*}+\xi)(u\star v)=(-1)^{|u|}\xi(u)\star v+(-1)^{|u|+|v|}u\star(\xi^{*}+\xi)(v),$
together with
$(-1)^{|u|+|v|}\,\omega(\xi(\xi+\xi^{*})(u),v)-\omega(b_{0},u\star v)=(-1)^{|u|+|v|}\,\omega(\xi^{2}(u),v)+\omega((\xi+\xi^{*})(\xi(u)),v).$
Since both $\xi\circ(\xi+\xi^{*})$ and $\xi^{2}$ are even, and $\omega$ is even, it follows that when $|u|+|v|=\bar{1}$, the equality reduces to the trivial identity $0=0$. Consequently, the only nontrivial case occurs when $|u|+|v|=\bar{0}$, which yields the relation:
$\omega(\xi(\xi+\xi^{*})(u),v)+\omega(u\star b_{0},v)=\omega(\xi^{2}(u),v)+\omega((\xi+\xi^{*})(\xi(u)),v).$
Since $\omega$ is nondegenerate, we deduce that
$\xi\circ\xi^{*}+\xi^{*}\circ\xi=-3\xi^{2}+{\mathrm{R}}^{\star}_{b_{0}}.$
The identity
$(\widetilde{\xi}\circ\widetilde{D}-\widetilde{D}\circ\widetilde{\xi})(u)=(-1)^{|d||u|}\big(\widetilde{\xi}^{2}-{\mathrm{R}}^{\diamond}_{\widetilde{b_{0}}}\big)(u)$
is equivalent to
$\xi\circ\xi^{*}+\xi^{*}\circ\xi=-3\xi^{2}+{\mathrm{R}}^{\star}_{b_{0}},\text{ and }\omega(b_{0},\xi(u))=(-1)^{|u|}\omega(b_{0},\xi(u)).$
Since both $b_{0}$ and $\omega$ are even, and $\xi$ is odd, we deduce that $(-1)^{|u|}\omega(b_{0},\xi(u))=\omega(b_{0},\xi(u))$.
In addition, the relation
$(1-(-1)^{|d|})\big(\widetilde{D}^{2}-\lambda\widetilde{D}-{\mathrm{L}}^{\diamond}_{b_{0}}\big)=0,$
is equivalent to
$(2\xi+\xi^{*})(b_{0})=0,\quad\text{and}\quad(\xi+\xi^{*})\circ(\xi+\xi^{*})=-{\mathrm{L}}^{\star}_{b_{0}}.$
Finally, the identity $(\widetilde{\xi}-\widetilde{D})(\widetilde{b_{0}})=0$ is equivalent to $(2\xi+\xi^{*})(b_{0})=0.$
Since $(V,\mathfrak{b},D,\xi^{*},b_{0},0)$ is a context of almost semi-direct product of left-symmetric superalgebras, it follows that $(V,\mathcal{H}:=\mathfrak{b}\oplus V^{*},\widetilde{D},\widetilde{\xi},\widetilde{b_{0}},0)$ is also a context of almost semi-direct product of left-symmetric superalgebras if and only if the conditions (29) are satisfied. This proves the lemma. ∎
Now, we are in a position to introduce the odd flat double extensions of flat orthosymplectic quasi-Frobenius Lie superalgebras.
Theorem.
Let $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ be a flat orthosymplectic quasi-Frobenius Lie superalgebra, and let $\star_{\mathfrak{b}}$ denote the natural symplectic product associated with it. Let $V=\mathbb{K}d$ be a purely odd one-dimensional vector superspace and $V^{*}=\mathbb{K}e$ its dual. Assume that there exist an odd homogeneous linear map $\xi:\mathfrak{b}\to\mathfrak{b}$, and an even element $b_{0}\in\mathfrak{b}_{0}$, satisfying the system (29). Define the vector superspace $\mathfrak{g}:=\mathbb{K}d\oplus\mathfrak{b}\oplus\mathbb{K}e,$ equipped with the Lie bracket and even bilinear form $\omega$ as follows:
(30) $\begin{array}[]{lcllcl}[d,u]&=&-(-1)^{|u|}(\xi^{*}+2\xi)(u)+\omega_{\mathfrak{b}}(b_{0},u)e,&[d,d]&=&2b_{0},\\[5.69054pt] [u,v]&=&[u,v]_{\mathfrak{b}}+\big((-1)^{|v|}\omega_{\mathfrak{b}}(\xi(u),v)+(-1)^{|u|}\omega_{\mathfrak{b}}(\xi^{*}(u),v)\big)e,\end{array}$
for all $u,v\in\mathfrak{b}$ and
$\omega|_{\mathfrak{b}\times\mathfrak{b}}=\omega_{\mathfrak{b}},\qquad\omega(e,d)=\omega(d,e)=1,\qquad\omega(d,\mathfrak{b})=\omega(e,\mathfrak{b})=0.$
Then, $(\mathfrak{g},[\;,\;],\omega)$ is a flat orthosymplectic quasi-Frobenius Lie superalgebra.
Furthermore, the natural symplectic product $\star$ on associated with $(\mathfrak{g},[\;,\;],\omega)$, is given by
(31) $\begin{array}[]{lll}e\star u=u\star e=0,&e\star d=d\star e=0,&d\star u=-(-1)^{|u|}(\xi^{*}+\xi)(u)+\omega_{\mathfrak{b}}(b_{0},u)e,\\[3.00003pt] u\star d=\xi(u),&d\star d=b_{0},&u\star v=u\star_{\mathfrak{b}}v+(-1)^{|v|}\omega_{\mathfrak{b}}(\xi(u),v)e,\end{array}$
for all $u,v\in\mathfrak{b}$.
The flat orthosymplectic quasi-Frobenius Lie superalgebra $(\mathfrak{g},[\;,\;],\omega)$ is called the odd flat double extension of $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ by means of $(\xi,b_{0})$.
Proof.
The proof follows the same lines as that of Theorem 5.5.5.  Theorem, using Lemmas 5.5.10.  Lemma and 5.5.11.  Lemma. ∎
Proposition.
Let $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ be a flat orthosymplectic quasi-Frobenius Lie superalgebra and let $(\mathfrak{g},[\;,\;],\omega)$ be the odd flat double extensionn of $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ by means of $(\xi,b_{0})$. If $(\mathfrak{b},[\;,\;]_{\mathfrak{b}})$ is nilpotent, then $(\mathfrak{g},[\;,\;])$ is nilpotent as well.
Proof.
Let us define the operator $\Delta$ by
$\Delta(u):=(-1)^{|u|}(2\xi+\xi^{*})(u).$
Then
$\Delta^{2}(u)=(-1)^{|u|+|d|}(2\xi+\xi^{*})\big((-1)^{|u|}(2\xi+\xi^{*})(u)\big)=-\big(4\xi^{2}+2\xi\circ\xi^{*}+2\xi^{*}\circ\xi+(\xi^{*})^{2}\big)(u).$
According to the system (29), we have
$\xi\circ\xi^{*}+\xi^{*}\circ\xi+3\xi^{2}={\mathrm{R}}^{\star_{\mathfrak{b}}}_{b_{0}},\text{ and }{\mathrm{L}}^{\star_{\mathfrak{b}}}_{b_{0}}=-(\xi^{2}+\xi\circ\xi^{*}+\xi^{*}\circ\xi+(\xi^{2})^{*}).$
It follows that $\Delta^{2}={\mathrm{L}}^{\star_{\mathfrak{b}}}_{b_{0}}-{\mathrm{R}}^{\star_{\mathfrak{b}}}_{b_{0}}=\mathop{\mathrm{ad}}\nolimits_{b_{0}}^{\mathfrak{b}}$. Since $(\mathfrak{b},[\,,\,]_{\mathfrak{b}})$ is nilpotent, it follows that $\Delta^{2}$ is nilpotent as well.
Now, if $(\mathfrak{g},[\;,\;],\omega)$ is an odd flat double extension of $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$, the Lie brackets in $\mathfrak{g}$ are given by
$d,u]=\Delta(u)+\omega_{\mathfrak{b}}(b_{0},u)e,\qquad[u,v]=[u,v]_{\mathfrak{b}}+\big((-1)^{|v|}\omega_{\mathfrak{b}}(\xi(u),v)+(-1)^{|u|}\omega_{\mathfrak{b}}(\xi^{*}(u),v)\big)e,$
for all $u,v\in\mathfrak{b}$. By induction, one can show that for any $k\in\mathbb{N}$, we have
$\mathop{\mathrm{ad}}\nolimits^{k}_{d}(u)=\Delta^{k}(u)+\omega_{\mathfrak{b}}(b_{0},D^{k-1}(u))e,$
and
$\mathop{\mathrm{ad}}\nolimits_{u}^{k}(v)=(\mathop{\mathrm{ad}}\nolimits_{u}^{\mathfrak{b}})^{k}(v)+\big((-1)^{\alpha}\omega_{\mathfrak{b}}(\xi(u),\mathop{\mathrm{ad}}\nolimits_{u}^{\mathfrak{b}})^{k-1}(v))+(-1)^{|u|}\omega_{\mathfrak{b}}(\xi^{*}(u),\mathop{\mathrm{ad}}\nolimits_{u}^{\mathfrak{b}})^{k-1}(v))\big)e,$
where $\alpha=|(\mathop{\mathrm{ad}}\nolimits_{u}^{\mathfrak{b}})^{k-1}(v)|$. Since $(\mathfrak{b},[\;,\;]_{\mathfrak{b}})$ is nilpotent, $\Delta$ is nilpotent and $e\in Z(\mathfrak{g})$, it follows that $(\mathfrak{g},[\;,\;])$ is nilpotent as well. This completes the proof. ∎
5.5.14. Even Flat double extension of flat periplectic Lie superalgebras
Following the approach of Theorem 5.5.5.  Theorem, we now define the even flat double extension for periplectic quasi-Frobenius Lie superalgebras. The proof, which relies on Lemmas 5.5.2.  Lemma and 5.5.3.  Lemma, yields the following result.
Theorem.
Let $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ be a flat periplectic quasi-Frobenius Lie superalgebra, and let $\star_{\mathfrak{b}}$ denote the natural symplectic product associated with it. Let $V=\mathbb{K}d$ be a one-dimensional vector space and and $\Pi(V^{*})=(\mathbb{K}e)_{\bar{1}}$, where $V^{*}$ its dual. Assume that there exist an even linear map $\xi:\mathfrak{b}\to\mathfrak{b}$, and an even element $b_{0}\in\mathfrak{b}_{\bar{0}}$, satisfying the system (20). Define the vector superspace $\mathfrak{g}:=V\oplus\mathfrak{b}\oplus\Pi(V^{*}),$ equipped with the Lie bracket and odd bilinear form $\omega$ as follows:
(32) $d,u]=(\xi^{*}-2\xi)(u)+\omega_{\mathfrak{b}}(b_{0},u)e,\quad\text{ and }\quad[u,v]=[u,v]_{\mathfrak{b}}+\omega_{\mathfrak{b}}\big((\xi+\xi^{*})(u),v\big)e,$
for all $u,v\in\mathfrak{b}$, and
$\omega|_{\mathfrak{b}\times\mathfrak{b}}=\omega_{\mathfrak{b}},\qquad\omega(e,d)=-\omega(d,e)=1,\qquad\omega(d,\mathfrak{b})=\omega(e,\mathfrak{b})=0.$
Then, $(\mathfrak{g},[\;,\;],\omega)$ is a flat periplectic quasi-Frobenius Lie superalgebra.
Furthermore, the natural symplectic product $\star$ associated with $(\mathfrak{g},[\;,\;],\omega)$, is given by
(33) $\begin{array}[]{lll}\displaystyle e\star u=u\star e=0,&e\star d=d\star e=0,&\displaystyle d\star u=(\xi^{*}-\xi)(u)+\tfrac{1}{3}\omega_{\mathfrak{b}}(b_{0},u)e,\\[3.00003pt] \displaystyle u\star d=\xi(u)-\tfrac{2}{3}\omega_{\mathfrak{b}}(b_{0},u)e,&d\star d=\tfrac{1}{3}b_{0},&u\star v=u\star_{\mathfrak{b}}v+\omega_{\mathfrak{b}}(\xi(u),v)e,\end{array}$
for all $u,v\in\mathfrak{b}$.
The flat periplectic quasi-Frobenius Lie superalgebra $(\mathfrak{g},[\;,\;],\omega)$ is called the even flat double extension of $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ by means of $(\xi,b_{0})$.
Proof.
The proof is similar to that of Theorem 5.5.5.  Theorem, using Lemmas 5.5.2.  Lemma and 5.5.3.  Lemma. ∎
5.5.16. Odd Flat double extension of flat periplectic Lie superalgebras
We now introduce the notion of odd flat double extension of flat periplectic quasi-Frobenius Lie superalgebras by a purely odd one-dimensional vector superspace.
Let $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ be a flat periplectic quasi-Frobenius Lie superalgebra, and let $\star_{\mathfrak{b}}$ denote the natural symplectic product associated with it. Let $V=\operatorname{\mathbb{K}}d$ be a purely odd one-dimensional vector superspace and $\Pi(V^{*})=\operatorname{\mathbb{K}}e$, where $V^{*}$ its dual.
To introduce the process of the odd flat double extension by $V$, we consider an odd linear map $\xi:\mathfrak{b}\rightarrow\mathfrak{b},$ and an even element $b_{0}\in\mathfrak{b}_{\bar{0}}$. We introduce the map $D:\mathfrak{b}\rightarrow\mathfrak{b}$ defined as
$D(u)=(-1)^{|u|}(\xi^{*}-\xi)(u).$
These two maps can be extended as $\widetilde{D},\widetilde{\xi}:\mathfrak{b}\oplus\Pi(V^{*})\rightarrow\mathfrak{b}\oplus\Pi(V^{*}),$ as follows:
(34) $\widetilde{D}(u+\alpha e)=(-1)^{|u|}(\xi^{*}-\xi)(u)+\,\omega(b_{0},u)\,e,\quad\widetilde{\xi}(u+\alpha e)=\xi(u),$
for all $u\in\mathfrak{b}$ and $\alpha\in\operatorname{\mathbb{K}}$. Put $\widetilde{b_{0}}=-b_{0}$.
Lemma.
The superspace $\mathcal{H}:=\mathfrak{b}\oplus\Pi(V^{*})$ equipped with the product
$(u+\alpha e)\diamond(v+\beta e):=u\star v+(-1)^{|v|}\omega(\xi(u),v)\,e,\qquad\forall\,u,v\in\mathfrak{b},\;\alpha,\beta\in\operatorname{\mathbb{K}},$
is a left-symmetric superalgebra if and only if
(35) $\xi([u,v])=\mathrm{L}_{u}^{\star}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\star}(\xi(u)),$
for all $u,v\in\mathfrak{b}$.
Proof.
The proof is similar to that of Lemma 5.5.2.  Lemma. ∎
Now, let us also assume that $(V,\mathfrak{b},D,\xi,-b_{0},0)$ is a context of almost semi-direct product of left-symmetric superalgebras, where $D(u)=(-1)^{|u|}(\xi^{*}-\xi)(u)$, for all $u\in\mathfrak{b}$.
Lemma.
$(V,\mathcal{H}:=\mathfrak{b}\oplus\Pi(V^{*}),\widetilde{D},\widetilde{\xi},\widetilde{b_{0}},0)$ is a context of almost semi-direct product of left-symmetric superalgebras if and only if
(36) $\left\{\begin{array}[]{rclrcl}\xi([u,v])&=&\mathrm{L}_{u}^{\star}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\star}(\xi(u)),&[\xi,\xi^{*}]&=&3\xi^{2}+{\mathrm{R}}^{\star}_{b_{0}},\\[5.69054pt] D(u\star v)&=&D(u)\star v+(-1)^{|u|}v\star D(v)+(-1)^{|u|}\xi(u)\star v,&{\mathrm{L}}^{\star}_{b_{0}}&=&(\xi+\xi^{*})^{2},\\[5.69054pt] 2\xi(b_{0})&=&\xi^{*}(b_{0}),\end{array}\right$.
for all $u,v\in\mathfrak{b}$.
Proof.
Let $u,v\in\mathfrak{b}$ and $\alpha,\beta\in\mathbb{K}$. A straightforward computation shows that the equality
$\widetilde{\xi}([u,v])=\mathrm{L}_{u}^{\diamond}(\widetilde{\xi}(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\diamond}(\widetilde{\xi}(u))$
is equivalent to
$\xi([u,v])=\mathrm{L}_{u}^{\star}(\xi(v))-(-1)^{|u||v|}\mathrm{L}_{v}^{\star}(\xi(u)).$
Moreover, the relation
$\widetilde{D}(u)\diamond v-\widetilde{D}(u\diamond v)=(-1)^{|d||u|}\big(\widetilde{\xi}(u)\diamond v-u\diamond\widetilde{D}(v)\big)$
is equivalent to
$(-1)^{|u|}(\xi^{*}-\xi)(u)\star v-(-1)^{|u|+|v|}(\xi^{*}-\xi)(u\star v)=(-1)^{|u|}\xi(u)\star v-(-1)^{|u|+|v|}u\star(\xi^{*}-\xi)(v),$
together with
$(-1)^{|u|+|v|}\,\omega(\xi(\xi^{*}-\xi)(u),v)-\omega(b_{0},u\star v)=(-1)^{|u|+|v|}\,\omega(\xi^{2}(u),v)+\omega((\xi^{*}-\xi)(\xi(u)),v).$
Since both $\xi\circ(\xi^{*}-\xi)$ and $\xi^{2}$ are even, and $\omega$ is odd, it follows that when $|u|+|v|=\bar{0}$, the equality reduces to the trivial identity $0=0$. Consequently, the only nontrivial case occurs when $|u|+|v|=\bar{1}$, leading to the following relation:
$\omega(\xi(\xi^{*}-\xi)(u),v)+\omega(u\star b_{0},v)=-\omega(\xi^{2}(u),v)+\omega((\xi^{*}-\xi)(\xi(u)),v).$
Since $\omega$ is nondegenerate, we deduce that
$\xi\circ\xi^{*}+\xi^{*}\circ\xi=3\xi^{2}+{\mathrm{R}}^{\star}_{b_{0}}.$
The identity
$(\widetilde{\xi}\circ\widetilde{D}-\widetilde{D}\circ\widetilde{\xi})(u)=(-1)^{|d||u|}\big(\widetilde{\xi}^{2}-{\mathrm{R}}^{\diamond}_{\widetilde{b_{0}}}\big)(u),$
is equivalent to
$\xi\circ\xi^{*}+\xi^{*}\circ\xi=3\xi^{2}+{\mathrm{R}}^{\star}_{b_{0}},\text{ and }\omega(b_{0},\xi(u))=(-1)^{|u|}\omega(b_{0},\xi(u)).$
Since $b_{0}$ is even, $\omega$ is odd, and $\xi$ is odd, then the relation $\omega(b_{0},\xi(u))=(-1)^{|u|}\omega(b_{0},\xi(u))$ is therefore always satisfied. In addition, the relation
$(1-(-1)^{|d|})\big(\widetilde{D}^{2}-\lambda\widetilde{D}-{\mathrm{L}}^{\diamond}_{b_{0}}\big)=0,$
is equivalent to
$(2\xi-\xi^{*})(b_{0})=0,\quad\text{and}\quad(\xi^{*}-\xi)\circ(\xi^{*}-\xi)={\mathrm{L}}_{b_{0}}.$
Finally, the identity $(\widetilde{\xi}-\widetilde{D})(\widetilde{b_{0}})=0$ is equivalent to $(2\xi-\xi^{*})(b_{0})=0.$
Since $(V,\mathfrak{b},D,\xi^{*},b_{0},0)$ is a context of almost semi-direct product of left-symmetric superalgebras, it follows that $(V,\mathcal{H}:=\mathfrak{b}\oplus\Pi(V^{*}),\widetilde{D},\widetilde{\xi},\widetilde{b_{0}},0)$ is also a context of almost semi-direct product of left-symmetric superalgebras if and only if the conditions (29) are satisfied. This proves the Lemma. ∎
Theorem.
Let $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ be a flat periplectic quasi-Frobenius Lie superalgebra, and let $\star_{\mathfrak{b}}$ denote the natural symplectic product associated with it. Let $V=V_{\bar{1}}=(\mathbb{K}d)_{\bar{1}}$ be a purely odd one-dimensional vector superspace and $\Pi(V^{*})=\mathbb{K}e$ where $V^{*}$ its dual. Assume that there exist an odd homogeneous linear map $\xi:\mathfrak{b}\to\mathfrak{b}$, and an even element $b_{0}\in\mathfrak{b}_{0}$, satisfying the system (36). Define the vector superspace $\mathfrak{g}:=\mathbb{K}d\oplus\mathfrak{b}\oplus\mathbb{K}e,$ equipped with the Lie bracket and odd bilinear form $\omega$ given as follows:
(37) $\begin{array}[]{lcllcl}[d,u]&=&(-1)^{|u|}(\xi^{*}-2\xi)(u)+\omega_{\mathfrak{b}}(b_{0},u)e,&[d,d]&=&-2b_{0},\\[5.69054pt] [u,v]&=&[u,v]_{\mathfrak{b}}+\big((-1)^{|v|}\omega_{\mathfrak{b}}(\xi(u),v)+(-1)^{|u|}\omega_{\mathfrak{b}}(\xi^{*}(u),v)\big)e,\end{array}$
for all $u,v\in\mathfrak{b}$ and
$\omega|_{\mathfrak{b}\times\mathfrak{b}}=\omega_{\mathfrak{b}},\qquad\omega(e,d)=-\omega(d,e)=1,\qquad\omega(d,\mathfrak{b})=\omega(e,\mathfrak{b})=0.$
Then $(\mathfrak{g},[\;,\;],\omega)$ is a flat periplectic quasi-Frobenius Lie superalgebra.
Furthermore, the natural symplectic product associated with $(\mathfrak{g},[\;,\;],\omega)$, is given by
(38) $\begin{array}[]{lll}e\star u=u\star e=0,&e\star d=d\star e=0,&d\star u=(-1)^{|u|}(\xi^{*}-\xi)(u)+\omega_{\mathfrak{b}}(b_{0},u)e,\\[3.00003pt] u\star d=\xi(u),&d\star d=-b_{0},&u\star v=u\star_{\mathfrak{b}}v+(-1)^{|v|}\omega_{\mathfrak{b}}(\xi(u),v)e.\end{array}$
for all $u,v\in\mathfrak{b}$.
The flat periplectic quasi-Frobenius Lie superalgebra $(\mathfrak{g},[\;,\;],\omega)$ is called the odd flat double extension of $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ by means of $(\xi,b_{0})$.
Proof.
The structure of this proof is analogous to that of Theorem 5.5.5.  Theorem; the key difference is the use of Lemmas 5.5.17.  Lemma and 5.5.18.  Lemma instead of the lemmas used there. ∎
6. Flat quasi-Frobenius Lie superalgebras are nilpotent
In this section, we show that all non-abelian flat quasi-Frobenius Lie superalgebras have a degenerate center. Moreover, every such Lie superalgebra can be obtained via flat double extensions. In particular, every flat orthosymplectic quasi-Frobenius Lie superalgebra can be constructed by a sequence of flat double extensions starting from an an abelian Lie superalgebra, while flat periplectic quasi-Frobenius Lie superalgebras can be obtained by a sequence of flat double extensions starting from the trivial algebra $\{0\}$. Furthermore, we show that every flat orthosymplectic quasi-Frobenius Lie superalgebra is not only solvable but nilpotent
Theorem.
Let $(\mathfrak{g},[\;,\;],\omega)$ be a flat quasi-Frobenius Lie superalgebra. Then $Z(\mathfrak{g})\neq\{0\}$. Moreover, if $(\mathfrak{g},[\;,\;])$ is a non abelian Lie superalgebra, then $Z(\mathfrak{g})$ is degenerate.
Proof.
We distinguish two cases.
The case where $\omega$ is odd. If $(\mathfrak{g},[\;,\;])$ is abelian, then it is clear that $Z(\mathfrak{g})\neq 0.$ If $(\mathfrak{g},[\;,\;])$ is non-abelian then according to Theorem 4.5.  Theorem, $(\mathfrak{g},[\;,\;])$ is nilpotent. It follows that $Z(\mathfrak{g})\cap[\mathfrak{g},\mathfrak{g}]\neq\{0\},$ and since $Z(\mathfrak{g})\subseteq[\mathfrak{g},\mathfrak{g}]^{\bot}$, it follows that $Z(\mathfrak{g})\cap Z(\mathfrak{g})^{\bot}\neq\{0\}$ and we are done.
The case where $\omega$ is even. If $(\mathfrak{g},[\;,\;])$ is abelian, then it is clear that $Z(\mathfrak{g})\neq 0.$ Now, we assume that $(\mathfrak{g},[\;,\;])$ is non-abelian. We need the following two lemmas.
Lemma.
The following implication holds:
(39) $Z(\mathfrak{g})\cap[\mathfrak{g},\mathfrak{g}]=\{0\}\;\Longrightarrow\;Z([\mathfrak{g},\mathfrak{g}])\cap N_{\ell}(\mathfrak{g},\star)=\{0\}.$
Proof.
Suppose that $Z(\mathfrak{g})\cap[\mathfrak{g},\mathfrak{g}]=\{0\}$ and let $u\in Z([\mathfrak{g},\mathfrak{g}])\cap N_{\ell}(\mathfrak{g},\star)$. Then, for any $w\in\mathfrak{g}$ and any $v\in[\mathfrak{g},\mathfrak{g}]$, and since ${\mathrm{L}}^{\star}_{u}=0$ together with (14), we have $\mathop{\mathrm{ad}}\nolimits_{u}^{*}=\mathop{\mathrm{ad}}\nolimits_{u}$. Hence,
$0=\omega([u,v],w)=(-1)^{|u||v|}\omega(v,[u,w]).$
[…]
With this observation in mind, let us prove our result by contradiction. According to Prop. 4.4.  Theorem, $(\mathfrak{g},[\;,\;])$ is solvable and $Z(\mathfrak{g})\cap[\mathfrak{g},\mathfrak{g}]=\{0\}$. We consider the sequence of vector subspaces defined by
$C_{0}=[\mathfrak{g},\mathfrak{g}]\cap[\mathfrak{g},\mathfrak{g}]^{\perp},\qquad C_{k}=[[\mathfrak{g},\mathfrak{g}],C_{k-1}],\quad k\geq 1.$
By virtue of Prop.  4.6.  Proposition, we have $C_{0}\neq\{0\}$, and according to Prop. 4.7.  Proposition, $C_{k}\subset N_{\ell}(\mathfrak{g},\star)$ for any $k\geq 1$.
Lemma.
For any $k\geq 1$, if $C_{k}=\{0\}$ then $C_{k-1}=\{0\}$.
Proof.
Suppose that $C_{k}=\{0\}$. Then, for any $u\in C_{k-1}$, any $v\in[\mathfrak{g},\mathfrak{g}]$, and any $w\in\mathfrak{g}$, we have
$0=\omega([u,v],w)=\pm(-1)^{|u||v|}\omega(v,[u,w]),$
since $\mathop{\mathrm{ad}}\nolimits_{u}^{*}=-\mathop{\mathrm{ad}}\nolimits_{u}$ if $k=1$ and $\mathop{\mathrm{ad}}\nolimits_{u}^{*}=\mathop{\mathrm{ad}}\nolimits_{u}$ if $k\geq 2$. Thus, $u,w]\in[\mathfrak{g},\mathfrak{g}]^{\perp}\cap N_{\ell}(\mathfrak{g},\star)=Z(\mathfrak{g})$ and hence $u,w]\in Z(\mathfrak{g})\cap[\mathfrak{g},\mathfrak{g}]$ which shows that $u=0$. This completes the proof of the lemma. $\blacksquare$
Since $C_{0}\neq\{0\}$, we deduce that for all $k\geq 1$, $C_{k}\neq\{0\}$. But $\mathfrak{g},\mathfrak{g}]$ is nilpotent, and therefore there exists $k_{0}$ such that $C_{k_{0}}\subseteq Z([\mathfrak{g},\mathfrak{g}])$. We also have $C_{k_{0}}\subseteq N_{\ell}(\mathfrak{g},\star)$, according to (39), we obtain that $Z(\mathfrak{g})\cap[\mathfrak{g},\mathfrak{g}]\neq{0}$. Since $\mathfrak{g},\mathfrak{g}]\subseteq Z(\mathfrak{g})^{\perp}$, we deduce that $Z(\mathfrak{g})\cap[\mathfrak{g},\mathfrak{g}]\subseteq Z(\mathfrak{g})\cap Z(\mathfrak{g})^{\perp}\neq\{0\}$.∎
Remark.
Every flat quasi-Frobenius Lie superalgebra of dimension $2$ is abelian. Indeed, Theorem 6.1.  Theorem implies $Z(\mathfrak{g})\neq 0$. Hence, there exists a non-zero element $a\in Z(\mathfrak{g})_{\bar{i}}$.
If $\omega$ is odd, then $\dim(\mathfrak{g}_{\bar{0}})=\dim(\mathfrak{g}_{\bar{1}})$. Assume $a\in Z(\mathfrak{g})_{\bar{0}}$. We can find an element $b\in\mathfrak{g}_{\bar{1}}$ such that $\omega(a,b)=1$ and $b,b]=\alpha a$, where $\alpha\in\mathbb{K}$. As $\omega$ is closed, we obtain
$3\,\omega([b,b],b)=0\quad\Rightarrow\quad 3\alpha=0,$
which implies $\alpha=0$, and hence $(\mathfrak{g},[\;,\;])$ is abelian. The case where $a\in Z(\mathfrak{g})_{\bar{1}}$ is obvious.
[…]
Proposition.
Let $(\mathfrak{g},[\,,\,],\omega)$ be a flat quasi-Frobenius Lie superalgebra. Let $0\not=a\in Z(\mathfrak{g})\cap Z(\mathfrak{g})^{\perp}$, and denote $I:=\mathbb{K}a,$ and $\mathfrak{b}:=I^{\perp}/I.$ Let $\pi_{\mathfrak{b}}:I^{\perp}\to\mathfrak{b}$ be the canonical projection. We have
  1. $(i)$
    II and $I^{\perp}$ are totally isotropic graded two-sided ideals of $(\mathfrak{g},\star)$.
  2. $(ii)$
    The quotient $\mathfrak{b}=I^{\perp}/I$ admits a canonical structure of a flat quasi-Frobenius Lie superalgebra, defined by
    $\pi_{\mathfrak{b}}(u),\pi_{\mathfrak{b}}(v)]_{\mathfrak{b}}=\pi_{\mathfrak{b}}([u,v])\quad\text{and}\quad\omega_{\mathfrak{b}}(\pi_{\mathfrak{b}}(u),\pi_{\mathfrak{b}}(v))=\omega(u,v),$
    for all $u,v\in I^{\perp}$.
Proof.
Let us proof Part (i). Since $Z(\mathfrak{g})$ is a homogeneous ideal, the element $a$ can be chosen to be homogeneous. Since $a\in Z(\mathfrak{g})$, we have $a\star u=u\star a=0$ for all $u\in\mathfrak{g}$, which implies that $I$ is a totally isotropic graded two-sided ideal of $(\mathfrak{g},\star)$.
Now, let $v\in I^{\perp}$ and $w\in\mathfrak{g}$. We have
$\omega(w\star v,a)=(-1)^{|w||v|}\omega(v,w\star a)=0,\qquad\omega(v\star w,a)=(-1)^{|w||v|}\omega(w,v\star a)=0,$
hence $w\star v,\,v\star w\in I^{\perp}$.
If $\omega$ is even, then $I^{\perp}$ is a graded two-sided ideal of $(\mathfrak{g},\star)$.
If $\omega$ is odd, we distinguish two cases. If $a$ is even, and since $\omega$ is odd, we have $\mathfrak{g}_{\bar{0}}\subseteq I^{\perp}$, and therefore $(I^{\perp})_{\bar{0}}=\mathfrak{g}_{\bar{0}}$. If $a$ is odd, then $\mathfrak{g}_{\bar{1}}\subseteq I^{\perp}$, and hence $(I^{\perp})_{\bar{1}}=\mathfrak{g}_{\bar{1}}$. In both cases, we conclude that $I^{\perp}$ is also a graded two-sided ideal of $(\mathfrak{g},\star)$.
Since $I^{\perp}$ is a graded two-sided ideal of $(\mathfrak{g},\star)$, the quotient $\mathfrak{b}=I^{\perp}/I$ naturally inherits the structure of a Lie superalgebra and of a quasi-Frobenius. Indeed, for $u,v\in I^{\perp}$, we define
$\pi_{\mathfrak{b}}(u),\pi_{\mathfrak{b}}(v)]_{\mathfrak{b}}=\pi_{\mathfrak{b}}([u,v]),\qquad\omega_{\mathfrak{b}}(\pi_{\mathfrak{b}}(u),\pi_{\mathfrak{b}}(v))=\omega(u,v).$
These operations are well-defined since $I$ is totally isotropic, i.e., $\omega(u,a)=0$ for all $u\in I^{\perp}$ and $a\in I$.
Moreover, since $(\mathfrak{g},[\,,\,],\omega)$ is quasi-Frobenius then $(\mathfrak{b},[\,,\,]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ is quasi-Frobenius as well. Furthermore, the product defined on $\mathfrak{b}$ is given by
$\omega_{\mathfrak{b}}\left(\pi_{\mathfrak{b}}(u)\star_{\mathfrak{b}}\pi_{\mathfrak{b}}(v),\pi_{\mathfrak{b}}(w)\right)$ $=\frac{1}{3}\left(\omega_{\mathfrak{b}}\left([\pi_{\mathfrak{b}}(u),\pi_{\mathfrak{b}}(v)],\pi_{\mathfrak{b}}(w)\right)+\omega_{\mathfrak{b}}\left([\pi_{\mathfrak{b}}(u),\pi_{\mathfrak{b}}(w)],\pi_{\mathfrak{b}}(v)\right)\right)$
$=\frac{1}{3}\left(\omega\left([u,v],\omega\right)+\omega\left([u,w],v\right)\right)$
$=\omega(u\star v,w)$
Thus, we deduce that $(\mathfrak{b},\star_{\mathfrak{b}})$ is a left-symmetric superalgebra and hence $(\mathfrak{b},[\,,\,]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ is a flat quasi-Frobenius Lie superalgebra. ∎
Definition.
The Lie superalgebra $\mathfrak{b}$ obtained in Prop. 6.5.  Proposition is called the flat orthosymplectic quasi-Frobenius Lie superalgebra deduced from $I^{\perp}$ by means of $I$.
We now prove the converse of Theorems 5.5.5.  Theorem and 5.5.12.  Theorem.
Theorem.
Let $(\mathfrak{g},[\,,\,],\omega)$ be a flat orthosymplectic quasi-Frobenius non-abelian Lie superalgebra. Then $(\mathfrak{g},[\,,\,],\omega)$ is either:
  1. $(i)$
    an even flat double extension of a flat orthosymplectic quasi-Frobenius Lie superalgebra $(\mathfrak{b},[\,,\,]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ by means of $(\xi,b_{0})$, or
  2. $(ii)$
    an odd flat double extension of a flat orthosymplectic quasi-Frobenius Lie superalgebra $(\mathfrak{b},[\,,\,]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ by means of $(\xi,b_{0})$.
Proof.
Theorem 6.1.  Theorem implies $Z(\mathfrak{g})\cap Z(\mathfrak{g})^{\perp}\neq 0$. By Prop. 6.5.  Proposition, there exists a totally isotropic two-sided graded ideal $I=\operatorname{\mathbb{K}}e$ of $(\mathfrak{g},\star)$, and its orthogonal $I^{\perp}$ is also a two-sided graded ideal of $(\mathfrak{g},\star)$. We now distinguish two cases.
The case where $e$ is even. Since $\omega$ is even and non-degenerate, there exists an element $d\in\mathfrak{g}_{\bar{0}}\setminus\{0\}$ such that $\omega(e,d)=-\omega(d,e)=1.$
Since $I\subseteq I^{\perp}$, there exists a subsuperspace $\mathfrak{b}$ such that $I^{\perp}=I\oplus\mathfrak{B}$ and the restriction $\omega_{\mathfrak{B}}=\omega|_{\mathfrak{B}\times\mathfrak{B}}$ is non-degenerate. Let us write $\mathfrak{g}=\mathbb{K}e\oplus{\mathfrak{B}}\oplus\mathbb{K}d,$ and since $I^{\perp}=I\oplus{\mathfrak{B}}$ is a two-sided graded ideal of $(\mathfrak{g},\star)$, for any elements $u,v\in{\mathfrak{B}}$, we have
$u\star v=u\star_{\mathfrak{B}}v+\mu(u,v)e,$
where $\mu:{\mathfrak{B}}\times{\mathfrak{B}}\to\operatorname{\mathbb{K}}$ is an even bilinear form, and $\star_{\mathfrak{B}}:{\mathfrak{B}}\times{\mathfrak{B}}\to{\mathfrak{B}}$ is an even bilinear map.
It follows that $({\mathfrak{B}},\star_{\mathfrak{B}})$ is a left-symmetric superalgebra, $\mu$ satisfies the compatibility condition (16), and the restriction $\omega_{\mathfrak{B}}=\omega|_{{\mathfrak{B}}\times{\mathfrak{B}}}$ is orthosymplectic and closed on $({\mathfrak{B}},[\;,\;]_{\mathfrak{B}})$. Moreover, the canonical projection $\pi:{\mathfrak{B}}\rightarrow I^{\perp}/I=\mathfrak{b}$, $\pi(u)=\pi_{\mathfrak{b}}(u)$, for all $u\in{\mathfrak{B}},$ is an isomorphism of left-symmetric superalgebras. Thus, we can identify ${\mathfrak{B}}\cong I^{\perp}/I=\mathfrak{b}$, and hence $\mathfrak{g}=\mathbb{K}e\oplus\mathfrak{b}\oplus\mathbb{K}d,$ where $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})=(I^{\perp}/I,[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ is a flat orthosymplectic quasi-Frobenius Lie superalgebra.
Since $I^{\perp}=\mathbb{K}e\oplus\mathfrak{b}$ is a two-sided graded ideal of $(\mathfrak{g},\star)$, it is also a graded ideal of $(\mathfrak{g},[\;,\;])$. Then, for all $u,v\in\mathfrak{b}$, the Lie brackets on $\mathfrak{g}$ are given by
$d,u]=D(u)+T(u)e,\qquad[u,v]=[u,v]_{\mathfrak{b}}+\big(\mu(u,v)-(-1)^{|u||v|}\mu(v,u)\big)e,$
where $D\in\mathrm{End}(\mathfrak{b})_{\bar{0}}$, and $T:\mathfrak{b}\to\operatorname{\mathbb{K}}$ is an even linear map.
Since $\omega$ is even and non-degenerate, there exists $b_{0}\in\mathfrak{b}_{\bar{0}}$ such that $T(u)=\omega_{\mathfrak{b}}(b_{0},u)$. Moreover, the product associated with $(\mathfrak{g},[\,,\,],\omega)$ is given by
$u\star v$ $=u\star_{\mathfrak{b}}v+\mu(u,v)e,\;$ $u\star d=\xi(u)+f(u)e,$
$d\star u$ $=\rho(u)+g(u)e,\;$ $d\star d=\lambda d+c_{0}+\alpha e,$
where $\xi,\rho\in\mathrm{End}(\mathfrak{b})_{\bar{0}}$, $f,g:\mathfrak{b}\to\operatorname{\mathbb{K}}$ are even linear maps, and $\alpha,\lambda\in\operatorname{\mathbb{K}}$.
Using the natural symplectic product (13), for any $u,v\in\mathfrak{b}$ we have
$\mu(u,v)=\omega(u\star v,d)=-(-1)^{|u||v|}\omega(v,u\star d)=\omega_{\mathfrak{b}}(\xi(u),v),$
and hence
$\omega_{\mathfrak{b}}(\xi(u),v)=\omega(u\star v,d)=\tfrac{1}{3}([u,v],d)+\tfrac{1}{3}\omega([u,d],v)=\tfrac{1}{3}\omega_{\mathfrak{b}}((\xi+\xi^{*}-D)(u),v),$
which gives $D=\xi^{*}-2\xi$. Similarly,
$g(u)=\omega(d\star u,d)=\tfrac{1}{3}\omega_{\mathfrak{b}}(b_{0},u),$
and since $d,u]=d\star u-u\star d$, we obtain $\rho(u)=(\xi^{*}-\xi)(u)$ and $f(u)=-\tfrac{2}{3}\omega_{\mathfrak{b}}(b_{0},u)$.
Moreover,
$\omega(d\star d,e)=-\omega(d,d\star e)=\lambda=0,\quad\omega(d\star d,d)=\alpha=\omega([d,d],d)=0,$
and
$\omega_{\mathfrak{b}}(c_{0},u)=\omega(d\star d,u)=-\omega(d,d\star u)=\tfrac{1}{3}\omega_{\mathfrak{b}}(b_{0},u),$
hence $c_{0}=\tfrac{1}{3}b_{0}$.
Therefore, the Lie bracket and the associated product take the simplified form:
$d,u]=(\xi^{*}-2\xi)(u)+\omega_{\mathfrak{b}}(b_{0},u)e,\qquad[u,v]=[u,v]_{\mathfrak{b}}+\omega_{\mathfrak{b}}((\xi+\xi^{*})(u),v)e,$
and
$u\star v$ $=u\star_{\mathfrak{b}}v+\omega_{\mathfrak{b}}(\xi(u),v)e,\;\quad\quad\quad u\star d=\xi(u)-\tfrac{2}{3}\omega_{\mathfrak{b}}(b_{0},u)e,$
$d\star u$ $=(\xi^{*}-\xi)(u)+\tfrac{1}{3}\omega_{\mathfrak{b}}(b_{0},u)e,\quad\;\;d\star d=\tfrac{1}{3}b_{0}.$
Now, it is easy to show that $(\mathfrak{g},\star)$ is a left-symmetric superalgebra if and only if the pair $(\xi,b_{0})$ satisfies the system (20). Hence, $(\mathfrak{g},[\,,\,],\omega)$ is the even flat double extension of the flat orthosymplectic quasi-Frobenius Lie superalgebra $(\mathfrak{b},[\,,\,]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ by means of $(\xi,b_{0})$.
The case, where $e$ is odd. Since $\omega$ is even and non-degenerate, there exists an element $d\in{\mathfrak{b}}_{\bar{1}}\setminus\{0\}$ such that $\omega(e,d)=\omega(d,e)=1.$
Since $I\subseteq I^{\perp}$, there exists a subsuperspace $\mathfrak{b}$ such that $I^{\perp}=I\oplus\mathfrak{B}$ and the restriction $\omega_{\mathfrak{B}}=\omega|_{\mathfrak{B}\times\mathfrak{B}}$ is non-degenerate. Let us write $\mathfrak{g}=\mathbb{K}e\oplus{\mathfrak{B}}\oplus\mathbb{K}d,$ and since $I^{\perp}=I\oplus{\mathfrak{B}}$ is a two-sided graded ideal of $(\mathfrak{g},\star)$, for any elements $u,v\in{\mathfrak{B}}$, we have
$u\star v=u\star_{\mathfrak{B}}v+\mu(u,v)e,$
where $\mu:{\mathfrak{B}}\times{\mathfrak{B}}\to\operatorname{\mathbb{K}}$ is an even bilinear form, and $\star_{\mathfrak{B}}:{\mathfrak{B}}\times{\mathfrak{B}}\to{\mathfrak{B}}$ is an even bilinear map.
It follows that $({\mathfrak{B}},\star_{\mathfrak{B}})$ is a left-symmetric superalgebra, $\mu$ satisfies the compatibility condition (16), and the restriction $\omega_{\mathfrak{B}}=\omega|_{{\mathfrak{B}}\times{\mathfrak{B}}}$ is orthosymplectic and closed on $({\mathfrak{B}},[\;,\;]_{\mathfrak{B}})$. Moreover, the canonical projection $\pi:{\mathfrak{B}}\rightarrow I^{\perp}/I=\mathfrak{b}$, $\pi(u)=\pi_{\mathfrak{b}}(u)$, for all $u\in{\mathfrak{B}},$ is an isomorphism of left-symmetric superalgebras. Thus, we can identify ${\mathfrak{B}}\cong I^{\perp}/I=\mathfrak{b}$, and hence $\mathfrak{g}=\mathbb{K}e\oplus\mathfrak{b}\oplus\mathbb{K}d,$ where $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})=(I^{\perp}/I,[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ is a flat orthosymplectic quasi-Frobenius Lie superalgebra.
Since $I^{\perp}=\mathbb{K}e\oplus\mathfrak{b}$ is a two-sided graded ideal of $(\mathfrak{g},\star)$, it is also a graded ideal of $(\mathfrak{g},[\;,\;])$. Then, for all $u,v\in\mathfrak{b}$, the Lie brackets on $\mathfrak{g}$ are given by
$d,d]=c_{0},\quad[d,u]=D(u)+T(u)e,\qquad[u,v]=[u,v]_{B}+\big(\mu(u,v)-(-1)^{|u||v|}\mu(v,u)\big)e,$
where $D\in\mathrm{End}(\mathfrak{b})_{\bar{1}}$, and $T:\mathfrak{b}\to\operatorname{\mathbb{K}}$ is an even linear map.
Since $\omega$ is odd and non-degenerate, there exists $b_{0}\in\mathfrak{b}_{\bar{0}}$ such that $T(u)=\omega_{\mathfrak{b}}(b_{0},u)$. Moreover, the product associated with $(\mathfrak{g},[\,,\,],\omega)$ is given by
$u\star v$ $=u\star_{\mathfrak{b}}v+\mu(u,v)e,\;u\star d=\xi(u)+f(u)e,$
$d\star u$ $=\rho(u)+g(u)e,\quad\quad d\star d=a_{0},$
where $\xi,\rho\in\mathrm{End}(\mathfrak{b})_{\bar{1}}$, and $f,g:\mathfrak{b}\to\operatorname{\mathbb{K}}$ are even linear maps.
Using the natural symplectic product (13), for any $u,v\in\mathfrak{b}$, we have
$\mu(u,v)=\omega(u\star v,d)=-(-1)^{|u||v|}\omega(v,u\star d)=(-1)^{|v|}\omega_{\mathfrak{b}}(\xi(u),v),$
and hence
$(-1)^{|v|}\,\omega_{\mathfrak{b}}(\xi(u),v)$ $=\omega(u\star v,d)=\tfrac{1}{3}\omega([u,v],w)+\tfrac{1}{3}(-1)^{|v|}\omega([u,d],v)$
$=\tfrac{1}{3}(-1)^{|v|}\omega_{\mathfrak{b}}(\xi(u),v)+\tfrac{1}{3}(-1)^{|u|}\omega_{\mathfrak{b}}(\xi^{*}(u),v)-\tfrac{1}{3}(-1)^{|u|+|v|}\omega_{\mathfrak{b}}(D(u),v).$
Hence,
$\omega_{\mathfrak{b}}(D(u),v)=-2(-1)^{|u|}\omega_{\mathfrak{b}}(\xi(u),v)+(-1)^{|u|+|v|}\omega_{\mathfrak{b}}(\xi^{*}(u),v).$
Since both $\xi$ and $D$ are odd, and $\omega$ is even, it follows that when $|u|+|v|=\bar{0}$, the equality reduces to the trivial identity $0=0$. Consequently, the nontrivial case occurs when $|u|+|v|=\bar{1}$, leading to the following relation
$\omega_{\mathfrak{b}}(D(u),v)=-(-1)^{|u|}\omega_{\mathfrak{b}}((2\xi+\xi^{*})(u),v).$
It follows that $D(u)=-(-1)^{|u|}(2\xi+\xi^{*})(u)$. Similarly,
$f(u)=\omega(u\star d,d)=0,$
and since $d,u]=d\star u-u\star d$, we obtain $g(u)=\omega_{\mathfrak{b}}(b_{0},u)$ and $\rho(u)=(-1)^{|u|}(\xi^{*}+\xi)(u)$.
Moreover,
$\omega_{\mathfrak{b}}(a_{0},u)=\omega(d\star d,u)=\omega(d,d\star u)=\omega_{\mathfrak{b}}(b_{0},u),$
hence $a_{0}=b_{0}$ and since $d,d]=2d\star d=c_{0}=2b_{0}$.
Therefore, the Lie bracket and the associated product take the simplified form:
$\begin{array}[]{lcllcl}[d,u]&=&-(-1)^{|u|}(2\xi+\xi^{*})(u)+\omega_{\mathfrak{b}}(b_{0},u)e,&[d,d]&=&2b_{0},\\[5.69054pt] [u,v]&=&[u,v]_{\mathfrak{b}}+((-1)^{|v|}\omega_{\mathfrak{b}}(\xi(u),v)+(-1)^{|u|}\omega_{\mathfrak{b}}(\xi^{*}(u),v))e,\end{array}$
and
$u\star v$ $=u\star_{\mathfrak{b}}v+(-1)^{|v|}\omega_{\mathfrak{b}}(\xi(u),v)e,\quad\quad u\star d=\xi(u),$
$d\star u$ $=-(-1)^{|u|}(\xi^{*}+\xi)(u)+\omega_{\mathfrak{b}}(b_{0},u)e,\;d\star d=b_{0}.$
Now, it is easy to show that $(\mathfrak{g},\star)$ is a left-symmetric superalgebra if and only if the pair $(\xi,b_{0})$ satisfies the system (29). Hence, $(\mathfrak{g},[\,,\,],\omega)$ is the odd flat double extension of the flat orthosymplectic quasi-Frobenius Lie superalgebra $(\mathfrak{b},[\,,\,]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ by means of $(\xi,b_{0})$.∎
Corollary.
An orthosymplectic quasi-Frobenius Lie superalgebra $(\mathfrak{g},[\;,\;],\omega)$ is flat if and only if it can be obtained by a sequence of even or odd flat double extensions of flat orthosymplectic quasi-Frobenius Lie superalgebras, starting from the abelian superalgebra.
Proof.
If $(\mathfrak{g},[\;,\;])$ is non-abelian Lie superalgebra, then according to Theorem 6.7.  Theorem, $(\mathfrak{g},[\;,\;])$ is even or odd flat double extension of a flat orthosymplectic quasi-Frobenius Lie superalgebra $(\mathfrak{b}_{1},[\;,\;]_{\mathfrak{b}_{1}},\omega_{\mathfrak{b}_{1}})$ by means of $(\xi_{1},b_{1})$. If $(\mathfrak{b}_{1},[\;,\;]_{\mathfrak{b}_{1}})$ is a non-abelian Lie superalgebra, then it is itself an even or odd flat double extension of a flat orthosymplectic quasi-Frobenius Lie superalgebra $(\mathfrak{b}_{2},[\;,\;]_{\mathfrak{b}_{2}},\omega_{\mathfrak{b}_{2}})$ by means of $(\xi_{2},b_{2})$. This process has to stop since $\mathfrak{g}$ is finite-dimensional. ∎
The following theorem offers a sharpening of Theorem 4.4.  Theorem:
Theorem.
Every flat orthosymplectic quasi-Frobenius Lie superalgebra $(\mathfrak{g},[\;,\;],\omega)$ is nilpotent.
Proof.
We proceed by induction on $\dim\mathfrak{g}$. If $\dim\mathfrak{g}=1$ or $\dim\mathfrak{g}=2$, the result is trivial since in this case the superalgebra is abelian, and hence nilpotent (see Remark 6.4.  Remark).
Assume that the statement holds for every flat orthosymplectic quasi-Frobenius Lie superalgebra of total dimension $\leq n$, and let us prove it for dimension $n+1$.
Let $(\mathfrak{g},[\;,\;],\omega)$ be a flat orthosymplectic quasi-Frobenius Lie superalgebra of total dimension $n+1$. According to Theorem 6.7.  Theorem, $(\mathfrak{g},[\;,\;],\omega)$ is either an even or an odd flat double extension of a flat orthosymplectic quasi-Frobenius Lie superalgebra $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ of total dimension n1n-1.
By the induction hypothesis, $(\mathfrak{b},[\;,\;]_{\mathfrak{b}})$ is nilpotent. Then, by Props.  5.5.8.  Theorem and 5.5.13.  Proposition, we deduce that $(\mathfrak{g},[\;,\;],\omega)$ is nilpotent as well. ∎
Now, we give the converse of Theorem 5.5.15.  Theorem and Theorem 5.5.19.  Theorem.
Theorem.
Let $(\mathfrak{g},[\,,\,],\omega)$ be a flat periplectic quasi-Frobenius Lie superalgebra. Then $(\mathfrak{g},[\,,\,],\omega)$ is either:
  1. $(i)$
    an even flat double extension of a flat periplectic quasi-Frobenius Lie superalgebra $(\mathfrak{b},[\,,\,]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ by means of $(\xi,b_{0})$, or
  2. $(ii)$
    an odd flat double extension of a flat periplectic quasi-Frobenius Lie superalgebra $(\mathfrak{b},[\,,\,]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ by means of $(\xi,b_{0})$.
Proof.
The proof is similar to that of Theorem 6.7.  Theorem. ∎
Corollary.
A periplectic quasi-Frobenius Lie superalgebra $(\mathfrak{g},[\;,\;],\omega)$ is flat if and only if it can be obtained by a sequence of even or odd flat double extensions of flat periplectic quasi-Frobenius Lie superalgebras, starting from the trival superalgebra $\{0\}$.
Proof.
Theorem 6.9.1.  Theorem implies that $(\mathfrak{g},[\;,\;])$ is an even or odd flat double extension of a flat periplectic quasi-Frobenius Lie superalgebra $(\mathfrak{b}_{1},[\;,\;]_{\mathfrak{b}_{1}},\omega_{\mathfrak{b}_{1}})$ by means of $(\xi_{1},b_{1})$. The latter is itself an even or an odd flat double extension of a flat periplectic quasi-Frobenius Lie superalgebra $(\mathfrak{b}_{2},[\;,\;]_{\mathfrak{b}_{2}},\omega_{\mathfrak{b}_{2}})$ by means of $(\xi_{2},b_{2})$. Since the superdimension of $\mathfrak{g}$ is finite and even (see Lemma 2.1.  Lemma), there exists $k\in\mathbb{N}^{*}$ such that $(\mathfrak{b}_{k},[\;,\;]_{\mathfrak{b}_{k}})$ is trivial Lie superalgebra, i.e,. $\mathfrak{b}_{k}=\{0\}$. ∎
7. Classification of flat quasi-Frobenius Lie superalgebras of low dimensions
Here, we provide a classification of flat quasi-Frobenuis Lie superalgebras of dimension $\leq 5$. This classification is carried out over a field of characteristic zero that is algebraically closed. It is interesting to investigate the classification problem over any field.
7.1. Four-dimensional flat quasi-Frobenius Lie superalgebras
Here, we classify flat quasi-Frobenius Lie superalgebras of total dimension four. Theorems 6.7.  Theorem and 6.9.1.  Theorem imply that every four-dimensional flat quasi-Frobenius Lie superalgebra $(\mathfrak{g},[\;,\;],\omega)$ is an odd or even flat double extension of a two-dimensional abelian one $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$, defined by the pair $(\xi,b_{0})$. Indeed, every two-dimensional quasi-Frobenius Lie superalgebra is abelian (see Remark 6.4.  Remark).
Definition.
Let $(\mathfrak{b},[\,,]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ be a flat quasi-Frobenius Lie superalgebra. If $\omega_{\mathfrak{b}}$ is even, we say that a pair $(\xi,b_{0})\in\operatorname{End}(\mathfrak{b})\times\mathfrak{b}$ is even admissible (resp. odd admissible) if $(\xi,b_{0})$ satisfies System (20) (resp. System (29)). If $\omega_{\mathfrak{b}}$ is odd, we say that a pair $(\xi,b_{0})\in\operatorname{End}(\mathfrak{b})\times\mathfrak{b}$ is even admissible (resp. odd admissible) if $(\xi,b_{0})$ satisfies System (20) (resp. System (36)).
Remark.
Every flat orthosymplectic quasi-Frobenius Lie superalgebra of dimension three is abelian. Indeed, we suppose the Lie superalgbera is not abelian. Since $\omega$ is even, the even part $\mathfrak{g}_{0}$ has even dimension. It follows from Theorem 6.7.  Theorem that this $(\mathfrak{g},[\,,\,],\omega)$ is an even flat double extension of an one-dimensional flat orthosymplectic quasi-Frobenius Lie superalgebra $(\mathfrak{b},[\;,\;]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ by means of $(\xi,b_{0})$. Since $\omega_{\mathfrak{b}}$ is even, it followsthe $\mathfrak{b}_{\bar{0}}=0$. Moreover, as both $\xi$ and $b_{0}$ are even, it follows that $\xi=0$ and $b_{0}=0$. Hence, from (21), all brackets vanish, and therefore $(\mathfrak{g},[\,,\,],\omega)$ is abelian, a contradiction.
Proposition.
Let $(\mathfrak{b},[\,,\,]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ be the two-dimensional orthosymplectic quasi-Frobenius abelian Lie superalgebra.
  • (i)
    The pair $(\xi,b_{0})$ is even admissible if and only if there exists a basis $\mathbb{B}=\{e_{1},e_{2}\}$ of $\mathfrak{b}$ in which both elements are even (resp. both odd), such that the following conditions are satisfied:
    • If $\{e_{1},e_{2}\}$ are both even, then we either have
      $\xi=0,\quad b_{0}=\alpha e_{1}+\beta e_{2},\quad\omega=e_{1}^{*}\wedge e_{2}^{*},\quad\alpha,\beta\in\mathbb{K},\text{ or }$
      $\xi=\begin{pmatrix}0&a\\ 0&0\end{pmatrix},\quad b_{0}=\alpha e_{1},\quad\omega=e_{1}^{*}\wedge e_{2}^{*},$
      where $\alpha\in\mathbb{K}$ and $a\neq 0$.
    • If $\{e_{1},e_{2}\}$ are both odd, then we either have
      $\xi=0,\quad b_{0}=0,\quad\omega=e_{1}^{*}\wedge e_{1}^{*}+\epsilon\,e_{2}^{*}\wedge e_{2}^{*},\quad\epsilon=\pm 1.$
      or (where $a\neq 0$)
      $\xi=\begin{pmatrix}0&a\\ 0&0\end{pmatrix},\quad b_{0}=0,\quad\omega=-e_{1}^{*}\wedge e_{2}^{*}.$
  • (ii)
    The pair $(\xi,b_{0})$ is odd admissible if and only if there exists a basis $\mathbb{B}=\{e_{1},e_{2}\}$ of $\mathfrak{b}$ in which both elements are even (resp. both odd), such
    • If $\{e_{1},e_{2}\}$ are both even, we have
      $\xi=0,\quad b_{0}=\alpha e_{1}+\beta e_{2},\quad\omega=e_{1}^{*}\wedge e_{2}^{*},\quad\alpha,\beta\in\mathbb{K}.$
    • If $\{e_{1},e_{2}\}$ are both odd, we have
      $\xi=0,\quad b_{0}=0,\quad\omega=e_{1}^{*}\wedge e_{1}^{*}+\epsilon\,e_{2}^{*}\wedge e_{2}^{*},\quad\epsilon=\pm 1.$
Proof.
Let us only prove Part (i). If the pair $(\xi,b_{0})$ is even admissible, then
$\xi^{*}\circ\xi=0,\qquad\xi\circ\xi^{*}=\xi^{2},\qquad\text{and}\qquad b_{0}\in\ker(\xi^{*}-\xi).$
Hence, $\xi$ is nilpotent, and therefore $\xi=0$ or $\xi^{2}=0$ with $\xi\neq 0$. Since $\omega$ is even and $\mathfrak{b}$ is two-dimensional, we must have $\mathfrak{b}=\mathfrak{b}_{\bar{0}}$ or $\mathfrak{b}=\mathfrak{b}_{\bar{1}}$.
The case where $\mathfrak{b}=\mathfrak{b}_{\bar{0}}$, and $\xi\neq 0$. We can choose a symplectic basis $\{e_{1},e_{2}\}$ of $\mathfrak{b}$ such that
$\omega=e_{1}\wedge e_{2},\quad\xi=\begin{pmatrix}0&a\\ 0&0\end{pmatrix},\text{ and }b_{0}=\alpha e_{1}+\beta e_{2},$
where $a\neq 0$ and $\alpha,\beta\in\mathbb{K}$. The condition $(\xi^{*}-\xi)(b_{0})=0$ implies $\beta=0$.
The case where $\mathfrak{b}=\mathfrak{b}_{\bar{1}}$, and $\xi\neq 0$. We can choose a basis $\{f_{1},f_{2}\}$ of $(\mathfrak{b},\omega)$ such that
$\omega=-f_{1}^{*}\wedge f_{2}^{*},\quad\xi=\begin{pmatrix}0&a\\ 0&0\end{pmatrix},\text{ and }b_{0}=\alpha f_{1}+\beta f_{2},$
where $a\neq 0$ and $\alpha,\beta\in\mathbb{K}$. Again, $(\xi^{*}-\xi)(b_{0})=0$ implies $\beta=0$. If $\{f_{1},f_{2}\}$ are both odd, then $b_{0}$ being even forces $b_{0}=0$.
The case where $\xi=0$. The conclusion follows immediately. ∎
Proposition.
Let $(\mathfrak{b},[\,,\,]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ be the two-dimensional periplectic quasi-Frobenius abelian Lie superalgebra.
  • (i)
    The pair $(\xi,b_{0})$ is even admissible if and only if there exists a basis $\mathbb{B}=\{e_{1}\mid f_{1}\}$ of $\mathfrak{b}$ such that the following conditions are satisfied:
    $\xi=0,\;b_{0}=\alpha e_{1},\quad\text{and}\quad\omega=e_{1}^{*}\wedge f_{1}^{*},\quad\text{where}\quad\alpha\in\mathbb{K}.$
  • (ii)
    The pair $(\xi,b_{0})$ is odd admissible if and only if there exists a basis $\mathbb{B}=\{e_{1}\mid f_{1}\}$ of $\mathfrak{b}$ such that the following conditions are satisfied
    $\xi=0\quad b_{0}=\alpha e_{1}\quad\omega=e_{1}^{*}\wedge f_{1}^{*},\quad\text{where $\alpha\in\mathbb{K}$.}$
  • (iii)
    The pair $(\xi,b_{0})$ is odd admissible if and only if there exists a homogeneous basis $\mathbb{B}=\{e_{1}\mid f_{1}\}$ of $\mathfrak{b}$ such that $\omega=e_{1}^{*}\wedge f_{1}^{*}$, and one of the following holds:
    $\begin{array}[]{lll}\text{(1)}&\xi=0,&b_{0}=\alpha\,e_{1},\quad\alpha\in\mathbb{K},\\[4.0pt] \text{(2)}&\xi=\begin{pmatrix}0&0\\[2.0pt] a&0\end{pmatrix},&b_{0}=0,\qquad a\neq 0,\\[10.0pt] \text{(3)}&\xi=\begin{pmatrix}0&a\\[2.0pt] 0&0\end{pmatrix},&b_{0}=\alpha\,e_{1},\qquad a\neq 0,\;\alpha\in\mathbb{K},\end{array}$
    where $\alpha\in\mathbb{K}$.
Proof.
Let us only prove Part (i). If the pair $(\xi,b_{0})$ is even admissible, then
$\xi^{*}\circ\xi=0,\qquad\xi\circ\xi^{*}=\xi^{2},\qquad b_{0}\in\ker(\xi^{*}-\xi).$
Hence, $\xi$ is nilpotent. Since $\omega$ is odd and $\xi$ is even, there exists a homogeneous basis $\{e_{1}\mid f_{1}\}$ of $\mathfrak{b}$ such that
$\omega=e_{1}^{*}\wedge f_{1}^{*},\qquad\xi(e_{1})=a\,e_{1},\qquad\xi(f_{1})=b\,f_{1},\quad b_{0}=\alpha e_{1}$
for some scalars $a,b,\alpha\in\mathbb{K}$. Since $\xi$ is nilpotent, we must have $a=b=0$, and therefore $\xi=0$. ∎
Proposition.
Any four-dimensional flat quasi-Frobenius non-abelian Lie superalgebra is isomorphic to one of the following Lie superalgebras:
  • $(i)$
    $(\mathbb{K}\oplus\operatorname{\mathcal{H}}_{3},[\;,\;],\omega)$ (in the basis $\{x_{1},x_{2},x_{3}\}$):
    $x_{1},x_{2}]=x_{3},\quad\omega=x_{1}^{*}\wedge x_{4}^{*}+x_{2}^{*}\wedge x_{3}^{*},$
  • $(ii)$
    $(\mathfrak{g}^{2},[\;,\;],\omega)$ (in the basis $\{x_{1},x_{2}\mid y_{1},y_{2}\}$):
    $x_{1},y_{1}]=y_{2},\quad[y_{1},y_{1}]=x_{2}\mbox{ and }\omega=2x_{1}^{*}\wedge x_{2}^{*}-y_{1}^{*}\wedge y_{2}^{*},$
  • $(iii)$
    $(\mathfrak{g}^{3},[\;,\;],\omega)$ (in the basis $\{x_{1},x_{2}\mid y_{1},y_{2}\}$):
    $x_{1},y_{1}]=y_{2},\quad\omega=x_{1}^{*}\wedge y_{2}^{*}+x_{2}^{*}\wedge y_{1}^{*},$
  • $(iv)$
    $(\mathfrak{g}^{4},[\;,\;],\omega)$ (in the basis $\{x_{1},x_{2}\mid y_{1},y_{2}\}$):
    $y_{1},y_{1}]=x_{1},\quad[y_{1},y_{2}]=x_{2},\quad\omega=-2x_{1}^{*}\wedge y_{2}^{*}+x_{2}^{*}\wedge y_{1}^{*},$
Proof.
Let $(\mathfrak{g},[\,,\,],\omega)$ be a flat quasi-Frobenius non-abelian Lie superalgebra of total dimension $4$. Theorems 6.7.  Theorem and 6.9.1.  Theorem, implies that such a superalgebra is an even or an odd flat double extension of the two-dimensional abelian quasi-Frobenius Lie superalgebra $(\mathfrak{b},[\,,\,]_{\mathfrak{b}},\omega_{\mathfrak{b}})$ by means of $(\xi,b_{0})$. Following Prop. 7.1.3.  Proposition, we argue as follows.
The case where $\omega_{\mathfrak{b}}$ is even. We have several cases:
  • If $(\xi,b_{0})$ is an even admissible pair, we distinguish several cases:
    • If $\xi=0$, $b_{0}=\alpha e_{1}+\beta e_{2}$, and $\omega=e_{1}^{*}\wedge e^{*}_{2}$, where $\alpha,\beta\in\mathbb{K}$, then Eq. (21), implies that there exists a basis $\{e,d,e_{1},e_{2}\}$ of $\mathfrak{g}$ such that the Lie brackets are given by
      $\left[d,e_{1}\right]=-\beta e,\;\left[d,e_{2}\right]=\alpha e,\quad\omega=e^{*}\wedge d^{*}+e_{1}^{*}\wedge e_{2}^{*}.$
      Since $(\mathfrak{g},[\;,\;])$ is non-abelian, then $(\alpha,\beta)\neq(0,0)$. If for instance $\beta\neq 0$, then we choose
      $(x_{1},x_{2},x_{3},x_{4})=(-\frac{1}{\beta}e_{1},d,-e,-\beta e_{2}-\alpha e_{1}).$
      It follows that $(\mathfrak{g},\omega)$ is symplectomorphic to $(\mathbb{K}\oplus\operatorname{\mathcal{H}}_{3},[\;,\;],\omega)$.
    • If $\xi\neq 0$ and $b_{0}=\alpha e_{1}$, where $\alpha\in\mathbb{K}$, then Eq. (21), implies that there exists a basis $\{e,d,e_{1},e_{2}\}$ of $\mathfrak{g}$ such that the Lie brackets are given by
      $\left[d,e_{2}\right]=\alpha e-3ae_{1},\,\,\text{ where }a\neq 0.$
      Put
      $(x_{1},x_{2},x_{3},x_{4})=\left(3ad+\alpha e_{2},d+\frac{\alpha+1}{3a}e_{2},\alpha e-3ae_{1},-\frac{\alpha+1}{3a}e+e_{1}\right).$
      In this new basis, the Lie brackets and the form are reduced to
      $x_{1},x_{2}]=x_{3}\mbox{ and }\omega=x_{1}^{*}\wedge x_{4}^{*}+x_{2}^{*}\wedge x_{3}^{*},$
      which implies that $(\mathfrak{g},\omega)$ is symplectomorphic to $(\mathbb{K}\oplus\operatorname{\mathcal{H}}_{3},[\;,\;],\omega)$.
    • If $\xi\neq 0$ and $b_{0}=0$, then Eq. (21) implies that there exists a basis $\{e,d\mid e_{1},e_{2}\}$ of $\mathfrak{g}$ such that the Lie brackets are given by
      $\left[d,e_{2}\right]=-ae_{1},\quad[e_{2},e_{2}]=2ae,\,\,a\neq 0.$
      Put
      $(x_{1},x_{2}\mid y_{1},y_{2})=\left(-\frac{1}{a}d,2ae,-e_{2},-e_{1}\right).$
      In this new basis, the Lie brackets and the form are reduced to
      $x_{1},y_{1}]=y_{2},\quad[y_{1},y_{1}]=x_{2}\mbox{ and }\omega=2x_{1}^{*}\wedge x_{2}^{*}-y_{1}^{*}\wedge y_{2}^{*},$
      which implies that $(\mathfrak{g},\omega)$ is symplectomorphic to $(\mathfrak{g}^{2},[\;,\;],\omega)$.
  • If $(\xi,b_{0})$ is an odd admissible pair, it follows that $\xi=0$, $b_{0}=\alpha e_{1}+\beta e_{2}$, and $\omega=e_{1}^{*}\wedge e^{*}_{2}$, where $\alpha,\beta\in\mathbb{K}$. Eq (30)implies that there exists a basis $\{e_{1},e_{2}\mid e,d\}$ of $\mathfrak{g}$ such that the Lie brackets and the form are given by
    $d,d]=2\alpha e_{1}+2\beta e_{1},\;\;\left[d,e_{1}\right]=-\beta e,\;\left[d,e_{2}\right]=\alpha e,\quad\omega=-e^{*}\wedge d^{*}+e_{1}^{*}\wedge e_{2}^{*}.$
    Since $(\mathfrak{g},[\;,\;])$ is non-abelian, then $(\alpha,\beta)\neq(0,0)$. If for instance $\beta\neq 0$, then we put
    $(x_{1},x_{2}\mid y_{1},y_{2})=(\frac{1}{\beta}e_{1},2\beta e_{2}+2\alpha e_{1},d,e).$
    Thus $(\mathfrak{g},\omega)$ is symplectomorphic to $(\mathfrak{g}^{2},[\;,\;],\omega)$.
The case where $\omega_{\mathfrak{b}}$ is odd. We have several cases:
  • If $(\xi,b_{0})$ is an even admissible pair, then $\xi=0$ and $b_{0}=\alpha e_{1}$, and $\omega=e_{1}^{*}\wedge f^{*}_{1}$, where $\alpha\in\mathbb{K}$. Eq. (32) implies that there exists a basis $\{d,e_{1}\mid e,f_{1}\}$ of $\mathfrak{g}$ such that the Lie brackets are given by
    $\left[d,f_{1}\right]=\alpha e,\quad\omega=e^{*}\wedge d^{*}+e_{1}^{*}\wedge f_{1}^{*}.$
    Since $(\mathfrak{g},[\;,\;])$ is non-abelian, then $\alpha\neq 0$. We put
    $(x_{1},x_{2}\mid y_{1},y_{2})=(d,-\alpha e_{1},-\frac{1}{\alpha}f_{1},-e).$
    Thus, $(\mathfrak{g},\omega)$ is symplectomorphic to $(\mathfrak{g}^{3},[\;,\;],\omega)$.
  • If $(\xi,b_{0})$ is an odd admissible pair, we distinguish several cases:
    • If $\xi=0$, $b_{0}=\alpha e_{1}$, and $\omega=e_{1}^{*}\wedge e^{*}_{2}$, where $\alpha\in\mathbb{K}$, then Eq.(37), implies that there exists a basis $\{e,e_{1}\mid d,f_{1}\}$ of $\mathfrak{g}$ such that the Lie brackets are given by
      $\left[d,d\right]=-2\alpha e_{1},\quad\left[d,f_{1}\right]=\alpha e,\quad\omega=e^{*}\wedge d^{*}+e_{1}^{*}\wedge f_{1}^{*}.$
      Since $(\mathfrak{g},[\;,\;])$ is non-abelian, then $\alpha\neq 0$. We put
      $(x_{1},x_{2}\mid y_{1},y_{2})=(-2\alpha e_{1},e,d,\frac{1}{\alpha}f_{1},).$
      Thus $(\mathfrak{g},\omega)$ is symplectomorphic to $(\mathfrak{g}^{4},[\;,\;],\omega)$.
    • If $\xi\neq 0$, $b_{0}=0$, and $\omega=e_{1}^{*}\wedge e^{*}_{2}$, then Eq.(37), implies that there exists a basis $\{e,e_{1}\mid d,f_{1}\}$ of $\mathfrak{g}$ such that the Lie brackets are given by
      $\left[d,e_{1}\right]=af_{1},\quad\omega=e^{*}\wedge d^{*}+e_{1}^{*}\wedge f_{1}^{*}.$
      Put
      $(x_{1},x_{2}\mid y_{1},y_{2})=(e_{1},-ae,-\frac{1}{a}d,f_{1},).$
      Thus $(\mathfrak{g},\omega)$ is symplectomorphic to $(\mathfrak{g}^{3},[\;,\;],\omega)$.
    • If $\xi\neq 0$, $b_{0}=\alpha e_{1}$, and $\omega=e_{1}^{*}\wedge e^{*}_{2}$, where $\alpha\in\mathbb{K}$ then Eq.(37), implies that there exists a basis $\{e,e_{1}\mid d,f_{1}\}$ of $\mathfrak{g}$ such that the Lie brackets are given by
      $\left[d,d\right]=-2\alpha e,\quad\left[d,f_{1}\right]=ae_{1}+\alpha e_{0},\quad\left[f_{1},f_{1}\right]=-2ae\quad\omega=e^{*}\wedge d^{*}+e_{1}^{*}\wedge f_{1}^{*}.$
      If $\alpha=0$, then we put
      $(x_{1},x_{2}\mid y_{1},y_{2})=(-\frac{2}{a}e,ae_{1},\frac{1}{a}f_{1},ad).$
      But if $\alpha\neq 0$, put
      $(x_{1},x_{2}\mid y_{1},y_{2})=\left(-\lambda e_{1}-\mu e,\;-\lambda\,e_{1}+\mu\,e,\;\frac{1}{2}\big(\lambda^{-1}\,d-\mu^{-1}\,f_{1}\big),\;\lambda^{-1}\,d+\mu^{-1}\,f_{1}\right),$
      where $\lambda=\frac{\sqrt[3]{\alpha^{2}}}{\sqrt[3]{a}}$ and $\mu=\frac{\sqrt[3]{a^{2}}}{\sqrt[3]{\alpha}}$. Thus, $(\mathfrak{g},\omega)$ is symplectomorphic to $(\mathfrak{g}^{4},[\;,\;],\omega)$.
Remark.
The Lie superalgebras ${\mathfrak{g}}^{2}$, ${\mathfrak{g}}^{3}$, and ${\mathfrak{g}}^{4}$ correspond to $C^{3}+A$, $\mathbb{K}\oplus C^{3}$, and $(2A_{1,1}+2A)_{\frac{1}{2}}^{3}$ of [B], respectively. See also [BM, BR] for more details.
7.3. Five-dimensional flat quasi-Frobenius Lie superalgebras
Here, we classify flat quasi-Frobenius Lie superalgebras of total dimension five. Nilpotent Lie superalgebras of total dimension 5 over the reals were classified by Hegazi in [H].
Proposition.
Let $(\mathfrak{b},\omega_{\mathfrak{b}})$ be the three-dimensional orthosymplectic quasi-Frobenius abelian Lie superalgebra.
  • (i)
    The pair $(\xi,b_{0})$ is even admissible if and only if there exists a basis $\mathbb{B}=\{e_{1},e_{2}\mid f_{1}\}$ of $\mathfrak{b}$ such that the following conditions are satisfied
    • Either
      $\xi=0,\quad b_{0}=\alpha e_{1}+\beta e_{2},\quad\omega=e_{1}^{*}\wedge e_{2}^{*}+f_{1}^{*}\wedge f_{1}^{*},\quad\alpha,\beta\in\mathbb{K}.$
    • Or
      $\xi=\begin{pmatrix}0&a&0\\ 0&0&0\\ 0&0&0\end{pmatrix},\;b_{0}=\alpha e_{1},\,\omega=e_{1}^{*}\wedge e_{2}^{*}+f_{1}^{*}\wedge f_{1}^{*},\quad 0\not=a,\alpha\in\mathbb{K}.$
    1. (ii)
      The pair $(\xi,b_{0})$ is odd admissible if and only if there exists a basis $\mathbb{B}=\{e_{1},e_{2}\mid f_{1}\}$ of $\mathfrak{b}$ such that
      $\xi=0,\quad b_{0}=\alpha e_{1}+\beta e_{2},\quad\omega=e_{1}^{*}\wedge e_{2}^{*}+f_{1}^{*}\wedge f_{1}^{*},\quad\alpha,\beta\in\mathbb{K}.$
Proof.
We distinguish two cases.
The case where the pair $(\xi,b_{0})$ is even admissible. It follows that
$\xi^{*}\circ\xi=0,\qquad\xi\circ\xi^{*}=\xi^{2},\qquad\text{and}\qquad b_{0}\in\ker(\xi^{*}-\xi).$
Hence, $\xi$ is nilpotent, and therefore either $\xi=0$ or $\xi\neq 0$. Since $\omega_{\mathfrak{b}}$ is even, then $\dim(\mathfrak{b}_{\bar{1}})=1$. Because $\xi$ is even and nilpotent, we have $\xi|_{\mathfrak{b}_{\bar{1}}}=0$. Choose a basis $\{e_{1},e_{2}\mid f_{1}\}$ of $\mathfrak{g}$, such that
$\omega_{\mathfrak{b}}=e_{1}^{*}\wedge e_{2}^{*}+f_{1}^{*}\wedge f_{1}^{*}.$
Since $b_{0}$ is even, we can write $b_{0}=\alpha e_{1}+\beta e_{2}$, where $\alpha,\beta\in\mathbb{K}.$
Finally, since $\xi$ is even and nilpotent, its matrix with respect to $\{e_{1},e_{2}\mid f_{1}\}$ is
$\xi=\begin{pmatrix}0&a&0\\ 0&0&0\\ 0&0&0\end{pmatrix},\qquad a\neq 0.$
Since $b_{0}\in\ker(\xi^{*}-\xi)$, we deduce that $\beta=0$.
The case where the pair $(\xi,b_{0})$ is odd admissible. It follows that
$\xi^{*}\circ\xi=-\xi\circ\xi^{*},\;\xi^{2}=0,\qquad\text{and}\qquad b_{0}\in\ker(2\xi+\xi^{*}).$
Hence, $\xi$ is nilpotent, and therefore either $\xi=0$ or $\xi\neq 0$ with $\xi^{2}=0$. Since $\omega_{\mathfrak{b}}$ is even, then $\dim(\mathfrak{b}_{\bar{1}})=1$. If $\xi|_{\mathfrak{b}_{\bar{1}}}=0$, choose a basis $\{e_{1},e_{2}\mid f_{1}\}$ such that
$\omega_{\mathfrak{b}}=e_{1}^{*}\wedge e_{2}^{*}+f_{1}^{*}\wedge f_{1}^{*}.$
Since $b_{0}$ is even, we can write $b_{0}=\alpha e_{1}+\beta e_{2}$, $\alpha,\beta\in\mathbb{K}.$ Then, the matrix of $\xi$ with respect to $\{e_{1},e_{2}\mid f_{1}\}$ is
$\xi=\begin{pmatrix}0&0&0\\ 0&0&0\\ a&0&0\end{pmatrix},\qquad a\in\mathbb{K}.$
Since $\xi^{*}\circ\xi=-\xi\circ\xi^{*}$, it follows that $a=0$ and hence $\xi=0$.
Now, if $\xi|_{\mathfrak{b}_{\bar{1}}}\neq 0$, choose a basis $\{e_{1},e_{2}\mid f_{1}\}$ such that
$\omega_{\mathfrak{b}}=e_{1}^{*}\wedge e_{2}^{*}+f_{1}^{*}\wedge f_{1}^{*}.$
Since $b_{0}$ is even, we can write $b_{0}=\alpha e_{1}+\beta e_{2}$, $\alpha,\beta\in\mathbb{R}$.
Then, the matrix of $\xi$ with respect to $\{e_{1},e_{2}\mid f_{1}\}$ is
$\xi=\begin{pmatrix}0&0&a\\ 0&0&0\\ \alpha&\beta&0\end{pmatrix},\qquad a\neq 0,\quad\alpha,\beta\in\mathbb{R}.$
Since $\xi^{2}=0$, we deduce that $\alpha=\beta=0$. Moreover, from $\xi^{*}\circ\xi=-\xi\circ\xi^{*}$, it follows that $a=0$, and hence $\xi=0$, which is a contradiction. ∎
Proposition.
Any five-dimensional flat quasi-Frobenius non-abelian Lie superalgebra is isomorphic to the following:
  • $(i)$
    $(\mathfrak{g}^{1},[\;,\;],\omega)$ (in the basis $\{x_{1},x_{2},x_{3},x_{3}\,|\,y_{1}\}$):
    $x_{1},x_{2}]=x_{3},\quad\omega=x_{1}^{*}\wedge x_{4}^{*}+x_{2}^{*}\wedge x_{3}^{*}+y_{1}^{*}\wedge y_{1}^{*},$
  • $(ii)$
    $(\mathfrak{g}^{2},[\;,\;],\omega)$ (in the basis $\{x_{1},x_{2}\,|\,y_{1},y_{2},y_{3}\}$):
    $x_{1},y_{1}]=y_{2},\quad[y_{1},y_{1}]=x_{2}\mbox{ and }\omega=2x_{1}^{*}\wedge x_{2}^{*}-y_{1}^{*}\wedge y_{2}^{*}+y_{3}^{*}\wedge y_{3}^{*}.$
Proof.
Let $(\mathfrak{g},\omega)$ be a flat orthosymplectic quasi-Frobenius non-abelian Lie superalgebra of total dimension $5$. According to Theorem 6.7.  Theorem, such an algebra is an odd or even flat double extension of the three-dimensional abelian quasi-Frobenius Lie superalgebra $(\mathfrak{b},\omega_{\mathfrak{b}})$ by means of $(\xi,b_{0})$. By Prop. 7.3.1.  Proposition, we have:
The case where $(\xi,b_{0})$ is an even admissible pair. We have the following cases:
  • $(i)$
    If $\xi=0$ and $b_{0}=\alpha e_{1}+\beta e_{2}$, and $\omega=e_{1}^{*}\wedge e^{*}_{2}+f_{1}^{*}\wedge f^{*}_{1}$, where $\alpha,\beta\in\mathbb{R}$, then according to (21), there exists a basis $\{e,d,e_{1},e_{2}\mid f_{1}\}$ of $\mathfrak{g}$ such that the Lie brackets are given by
    $\left[d,e_{1}\right]=-\beta e,\;\left[d,e_{2}\right]=\alpha e,\quad\omega=e^{*}\wedge d^{*}+e_{1}^{*}\wedge e_{2}^{*}+f_{1}^{*}\wedge f^{*}_{1}.$
    Since $(\mathfrak{g},[\;,\;])$ is non-abelian, then $(\alpha,\beta)\neq(0,0)$. If for instance $\beta\neq 0$, then we put
    $(x_{1},x_{2},x_{3},x_{4}\mid y_{1})=(-\frac{1}{\beta}e_{1},d,-e,-\beta e_{2}-\alpha e_{1},f_{1}).$
    Thus $(\mathfrak{g},\omega)$ is symplectomorphic to $(\mathfrak{g}^{1},[\;,\;],\omega)$.
  • $(ii)$
    If $\xi\neq 0$ and $b_{0}=\alpha e_{1}$ where $\alpha\in\mathbb{K}$, then according to (21), there exists a basis $\{e,d,e_{1},e_{2}\mid f_{1}\}$ of $\mathfrak{g}$ such that the Lie brackets are given by
    $\left[d,e_{2}\right]=\alpha e-3ae_{1},\,\,a\neq 0.$
    Put
    $(x_{1},x_{2},x_{3},x_{4}\mid y_{5})=\left(3ad+\alpha e_{2},d+\frac{\alpha+1}{3a}e_{2},\alpha e-3ae_{1},-\frac{\alpha+1}{3a}e+e_{1},f_{1}\right).$
    Thus, in this new basis, the Lie brackets and the form are reduced to
    $x_{1},x_{2}]=x_{3}\mbox{ and }\omega=x_{1}^{*}\wedge x_{4}^{*}+x_{2}^{*}\wedge x_{3}^{*}+y_{1}^{*}\wedge y_{1}^{*},$
    which implies that $(\mathfrak{g},\omega)$ is symplectomorphic to $(\mathfrak{g}^{1},[\;,\;],\omega)$.
The case where $(\xi,b_{0})$ is an odd admissible pair. It follows that $\xi=0$ and $b_{0}=\alpha e_{1}+\beta e_{2}$, and $\omega=e_{1}^{*}\wedge e^{*}_{2}+f_{1}^{*}\wedge f^{*}_{1}$, where $\alpha,\beta\in\mathbb{K}$, then according to (30), there exists a basis $\{e_{1},e_{2}\mid e,d,f_{1}\}$ of $\mathfrak{g}$ such that the Lie brackets are given by
$d,d]=2\alpha e_{1}+2\beta e_{1},\;\;\left[d,e_{1}\right]=-\beta e,\;\left[d,e_{2}\right]=\alpha e,\quad\omega=-e^{*}\wedge d^{*}+e_{1}^{*}\wedge e_{2}^{*}+f_{1}^{*}\wedge f^{*}_{1}.$
Since $(\mathfrak{g},[\;,\;])$ is non-abelian, then $(\alpha,\beta)\neq(0,0)$. If for instance $\beta\neq 0$, then we put
$(x_{1},x_{2}\mid y_{1},y_{2},y_{3})=(\frac{1}{\beta}e_{1},2\beta e_{2}+2\alpha e_{1},d,e,f_{1}).$
Thus, $(\mathfrak{g},\omega)$ is symplectomorphic to $(\mathfrak{g}^{2},[\;,\;],\omega)$. ∎